### v1

# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

Abstract
This paper presents CoverUp, a novel system that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models (LLMs). CoverUp
iteratively improves coverage, interleaving coverage analysis with
dialogs with the LLM to focus its attention on as yet uncovered
lines and branches. The resulting test suites significantly improve
coverage over the current state of the art:
compared to CodaMosa, a hybrid LLM / search-based software
testing system, CoverUp substantially improves coverage across the board.
On a per-module basis, CoverUp achieves median line coverage
of 81% (vs. 62%), branch coverage of 53% (vs. 35%) and
line+branch coverage of 78% (vs. 55%).
We show that CoverUp’s iterative, coverage-guided approach is crucial to its effectiveness, contributing to nearly half of its successes.
††footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
Test generation tools aim to increase the coverage of the program under test.
By generating tests that cover a wider range of scenarios and execution paths,
they can help uncover bugs missed by manually written tests.
Unfortunately, current approaches often fail to achieve substantial coverage.
Pynguin exemplifies the search-based software testing (SBST)
approach [1].
Starting from randomly created test cases, SBST employs genetic
algorithms to mutate the tests in an attempt to increase coverage.
Unfortunately, its search process can get
stuck. CodaMosa addresses this problem by tracking
Pynguin’s progress [2]; when it concludes that the search process has
stalled, it prompts a large language model (LLM) for a test. It then
uses that test to re-seed the SBST, allowing it to resume progress.
While CodaMosa greatly improves Pynguin’s performance and constitutes
the previous state of the art, the resulting code coverage can still remain
relatively low (e.g., 34% overall branch coverage;
see Section IV for details).
This paper proposes CoverUp, a novel approach that drives LLM-based test
generation with coverage analysis. CoverUp incorporates detailed
coverage information into prompts customized to the current state
of the test suite, focusing the LLM on code that lacks coverage. Once
the LLM generates a set of tests, CoverUp executes them and
measures the resulting coverage. If the LLM-generated tests fail to
sufficiently improve coverage or fail to run altogether, CoverUp
continues the dialogue with the LLM, requesting changes to improve
coverage or fix the error.
In doing so, CoverUp in effect further refines and clarifies the
prompt, resulting in tests that significantly improve coverage.
To our knowledge, CoverUp is the first test generation system
to prompt an LLM for tests specifying the lines and branches lacking coverage
and that iteratively refines that prompt using coverage analysis.
We evaluate CoverUp using CodaMosa as a baseline, and on the
same benchmark applications originally used to evaluate it. Using the
state-of-the-art OpenAI’s GPT-4 Turbo LLM for both, we find that
CoverUp yields higher coverage in 61% of the benchmarks,
significantly increasing all coverage
metrics, whether computed across all code or aggregated over the
various benchmarks.
CoverUp increases median line coverage from 62% to 81%,
median branch coverage from 35% to 53% and
median line+branch coverage from 55% to 78%.
Across all code, CoverUp increases
line coverage from 54% to 61%, branch coverage from 34% to 43%,
and line+branch coverage from 49% to 57%. We show that CoverUp’s iterative, coverage-guided approach
is key to its effectiveness, accounting for roughly half of its successes.
This paper makes the following contributions:
It presents CoverUp, a novel system that drives the
generation of high-coverage Python regression tests via a combination
of coverage analysis and large-language models;
It conducts an empirical analysis of CoverUp, showing that it
significantly advances the state of the art.
II Related Work
Automated test generation is a well-established field of research.
Among the various proposed methods are
specification-based [3],
random [4, 5],
feedback-directed random [6],
symbolic execution guided random (“concolic”) [7, 8, 9],
search-based software testing (SBST) [10, 11, 1, 12, 13]
and transformer based [14] approaches.
The success of large language models on various tasks has motivated
their application to software engineering;
Wang et al. survey 102 recent papers using LLMs for software testing [15].
This section focuses on previous work most closely related to CoverUp.
TestPilot prompts an LLM to generate JavaScript unit tests based on the
function under test’s implementation, its documentation and usage snippets [16].
Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
refine the prompt in case the new tests do not improve coverage.
Bareiß et al. study the performance of the Codex LLM on
Java test case generation, among other code generation tasks [17].
Its prompts contain the signature of the method under test, an example of test generation,
and the body of the method; it discards any tests that do not compile.
By contrast, CoverUp’s prompts are based on code segments lacking coverage and
they explicitly request tests to improve it.
CoverUp also continues the chat with the LLM in case of build failure, failing tests,
or lack of coverage.
As Section IV-D shows, this iterative dialogue is
responsibe for almost 50% of CoverUp’s successful test generation.
Vikram et al. discuss prompting LLMs based on API documentation
to generate property-based tests for Python [18, 19].
CoverUp instead bases its prompting on the source code and
coverage measurements and, while it accepts property-based tests
if the LLM generates them, it does not explicitly request them.
TiCoder prompts LLMs to generate tests based on
a natural language description of the intended functionality of code [20].
Rather than facilitate regression testing, however, TiCoder
generates tests with the goal of clarifying and formalizing user
intent.
Fuzz4All [21] uses two separate LLMs to fuzz test programs in various programming
languages.
The distillation LLM takes in arbitrary user input, such as documentation
and code examples, and generates prompts for the generation LLM, which
generates test inputs.
After the initial user input distillation and prompt selection, Fuzz4All repeatedly
employs the generation LLM, mutating its prompt to produce additional test inputs.
While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches
differ fundamentally: Fuzz4All generates test inputs based on documentation or examples,
relying on an user-provided test oracle for testing; CoverUp instead works based on
the source code and coverage measurements, looking to add tests that improve the test
suite’s coverage.
CodaMosa [2], mentioned in the introduction, prompts
the Codex LLM for a test when its SBST algorithm gets stuck. Like
CoverUp, CodaMosa selects code to prompt for based on its (lack
of) coverage, and the prompt contains source code. However, it does
not use coverage information within the prompt itself, nor does it
systematically prompt for tests for all code lacking coverage; it only
prompts the LLM when necessary to help SBST make progress.
CodaMosa represents the previous state of the art in test case generation
for Python, and Section IV uses it as the baseline
for comparison to CoverUp.
Figure 1: Overview: The execution steps of CoverUp.
After measuring the test suite’s coverage, CoverUp prompts the
LLM so as to focus it on code segments that lack coverage.
It checks each new test, either accepting it if it increases coverage
or continuing the dialogue if it needs improvements.
A final check helps ensure that the suite works as a whole.
Given its incremental nature, once done CoverUp can be run again
to refocus the LLM on any segments missed in previous passes.
III Technique
CoverUp first measures the code coverage of the existing test suite
and splits up the source code into segments that require more testing
(Section III-A).
It then prompts the LLM for tests for each of these segments, using a chat
interface
(Section III-B).
As the LLM generates tests, CoverUp executes them, again measuring coverage.
If the test does not compile, fails or lacks coverage, it continues the chat,
pointing out any problems and requesting improvements
(Section III-C).
Finally, having processed all code segments, CoverUp checks the extended
test suite, looking for any integration problems
(Section III-D).
Figure 1 provides a graphical overview of these steps.
III-ACode Segmentation
CoverUp’s first step is to measure the code coverage of any pre-existing test suite;
it uses SlipCover [22], a recently introduced coverage analyzer that has
near-zero overhead, for that task.
It then uses the AST of each source file missing coverage to identify code
segments that can be used for prompting.
The goal is to be able to provide to the LLM, with each prompt, a code excerpt
that is intelligible, provides enough context, includes the lines or branches
lacking coverage, and is as short as possible.
Keeping code segments short is important for several different reasons.
Older OpenAI LLMs such as GPT-3.5 and GPT-4, as well as on-premises LLMs like
Meta’s LLaMA have small input windows of up to a few
thousand tokens, limiting the maximum size of the code segment.
OpenAI’s most recent GPT-4 Turbo models allow for inputs of up to 128K tokens,
but this limit applies not only to the initial prompt, but to the total number of tokens
in the entire sequence of prompts and responses in the case of a continued chat.
Finally, OpenAI charges at a per-token rate in both prompts and responses,
so that it literally pays to be succinct.
Each code segment typically includes an entire function or method and all its contents.
If a segment contains a method, CoverUp also includes a few lines of the original
class definition to provide the LLM with context, but omits other methods
and definitions; see Figure 2 for an example.
In an effort to enhance code comprehension, CoverUp retains all comments in a
segment, although removing them could make the segment more concise.
Concretely, to identify the code segments in a source file, CoverUp first
computes a set of “interesting” lines: these are lines that either lack
coverage, or that are the source or the destination in branches that lack coverage.
For each of these lines, it then looks in the AST for a class, function or method
object containing that line.
If the object found is a class and it spans more lines than a configurable limit,
CoverUp adds that class definition as segment context and recursively looks
for another, smaller object containing the line.
This limit is intended to guide CoverUp towards acceptable excerpt sizes,
facilitating its use with language models with more limited input window sizes.
Algorithm 1 describes this process in more detail.
Figure 2: Example of a code segment with a class method:CoverUp created a code segment for a class method,
omitting over 200 lines of source code originally
present between listing lines 1 and 2.
The original code is from the flutils package.
III-BInitial Prompting
CoverUp prompts the LLM for tests for every code segment it identifies;
Figure 3 shows an example of an initial prompt.
It has the following structure (circled numbers identify sections in the example):
a statement assigning the LLM the persona [23]
of an “expert Python test-driven developer”, intended to help guide it towards
high quality tests;
a sentence pointing out the code excerpt (segment), identifying what file
and module it comes from, and stating what lines or branches do not execute
when tested.
The portion specifying the lines and branches missing coverage is compressed
using line ranges, simplifying the prompt and reducing token usage.
a request for a pytest test;
a series of other requests, such as “make sure that the new test is correct”
and “include assertions” to steer the result towards usable tests;
a request that the response only include the new Python tests
to facilitate extracting it from the response and to reduce token usage; and
the code segment with any relevant lines highlighted by prefixing
them with their line numbers.
Relevant lines are those that lack coverage or that are
part of a branch that lacks coverage.
23456Figure 3: Example of an initial prompt:
The initial prompt selects a persona (1),
identifies where the code comes from and states how it lacks coverage when tested (2),
asks for tests (3-5),
and shows the code segment (6).
Any lines relevant to the request, i.e., used to identify
missing coverage, are highlighted by prefixing them with their line numbers
(as in “19:” in part (6)).
The code shown is from the flutils package.
Figure 4: Excerpt of test calling to pytest.main:
We found that certain tests generated by GPT-4
included top-level, unconditional calls into pytest.main.
Such calls, executed when pytest loads the test file,
caused it to restart the search for all tests in the suite,
slowing it down immensely.
Part (4) of the prompt now requests omitting such calls.
We found that tests generated by GPT-4 often included top-level code calling
into pytest.main or into parts of the test itself.
Figure 4 shows an example.
While such top-level code can make sense in a standalone test file, Python
executes it as part of the loading process and doing so may greatly
disrupt pytest’s operation.
In fact, the calls in some of these early tests caused pytest to restart
its test discovery, slowing it down to the point that it became unusable.
For that reason, part 4 of the prompt includes instructions that the test not
include such calls.
Prompt part 4 also instructs that the tests “clean up […], so as not
to affect other tests”.
It also suggests the use of pytest-mock, a library useful for isolating
the software under test from other parts of the code, facilitating test cleanup.
Despite these instructions, we still observe test generations that include side effects;
Section III-D describes how CoverUp detects and handles
such tests.
As Figure 3 shows, in part 6 of the prompt, CoverUp
prefixes certain lines with their numbers, such as in 20: before line 20.
It adds such prefixes to all lines lacking coverage as well as to all lines
that are part of a branch that lacks coverage.
This improves the LLM’s understanding of the missing coverage:
we found that prompts that only indicated the line number at the beginning of the
excerpt led to more tests being generated that did not improve coverage.
If a code segment has no prior coverage, that is, every one of its lines lacks coverage,
then CoverUp uses a slightly modified initial prompt where part 2 simply
states “the code does not execute”.
Typically, the LLM will not generate enough tests to immediately yield high coverage
in response to this simpler prompt, but this can be improved upon with a subsequent
execution of CoverUp, where it refocuses the LLM on those lines missed in the first
iteration.
III-CVerification and Continued Chat
Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
again measuring coverage using SlipCover.
This step is made practical by SlipCover’s near-zero overhead:
using coverage.py, the only other alternative tool for Python,
introduces a median of 180% overhead [22].
If the new tests pass and increase coverage, CoverUp saves them.
But, if they do not increase coverage, or if they result in failures or errors,
CoverUp continues the chat session, pointing out the problem(s) and requesting
improvements.
Figures 5 and 6 show examples
of those prompts.
12Figure 5: Example of a coverage follow-up prompt:CoverUp indicates to the LLM that a line and
some branches still weren’t covered (1), asking that it correct the test (2).
123Figure 6: Example of an error follow-up prompt:CoverUp indicates to the LLM that an error occurred (1),
requests a repair (2) and includes an excerpt of the error messages (3).
In the original execution from which this example is taken,
the LLM responds with an usable test (not shown).
Prior to executing a new test, CoverUp looks for any Python modules used by
the test that are absent from the system.
These are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package, but could also be indicative of
a module required by the program under test that has not yet been installed.
CoverUp offers options to record the missing modules in a Python requirements.txt
format to facilitate making them available for a subsequent run.
III-DIntegration Check
While CoverUp only saves tests that pass when ran by themselves, it is still
possible that some of these do not clean up properly, or otherwise have
side effects (“state pollution”) that cause other tests to fail [24].
This issue is not unique to LLM-based test generation: Fraser et al. [25]
report side effects on 50% of test classes generated by EvoSuite [11],
an SBST test generator.
Figure 7 shows an example where a generated test
not only modifies, but also, in a misguided attempt to clean up, deletes global
symbols used by the code under test.
To attempt to detect and handle such tests, once CoverUp has processed all
outstanding code segments, it executes the entire suite in a single session.
If tests then fail, CoverUp attempts to identify the cause, disabling tests as
necessary by renaming their files so that pytest ignores them.
Such tests remain available for review and possible reactivation by the
user, as they are known to contribute to coverage.
CoverUp supports two different methods for handling failing tests.
It can simply disable all failing tests without attempting to ascertain
the cause of the errors.
This method is fast and allows the user to move on quickly to prompting
for more tests, but may lead to a significant loss in coverage as various
tests are disabled unnecessarily.
Alternatively, CoverUp can look for the test or tests actually causing the
failures by successively running subsets of the entire suite, an approach
adapted from Delta Debugging [26].
This method requires additional test runs, and thus more time, but is likely
to reduce the loss in coverage resulting from disabled tests.
Figure 7: Excerpt of a test with side effects:
This LLM-generated test overwrites global constants and then,
in a misguided attempt to clean up, deletes them.
The test succeeds if executed by itself, but when executed along with
other tests, the missing constants cause other tests to fail.
CoverUp detects these failures during its integration check and
optionally either disables all failing tests, or searches for the
test causing the failures.
III-EHandling Flaky Tests
Flaky tests are those that can both pass or fail inconsistently,
without any changes to the code under test [27, 28].
Their unreliable behavior has a variety of causes;
in the context of LLM generated tests, they can also in part
result from not providing the LLM sufficient information.
Figure 8 shows a test for the
mimesis data generator package.
The code under test randomly assigns a blood type to its
Person object; the LLM, not having been provided with
the list of valid names, assumes that it includes 0-
and 0+, when the actually code uses O-
and O+ (using the letter O, not zero).
Consequently, the test fails whenever that blood type is assigned.
Figure 8: Example of a flaky test:
This test checks that the Person’s
blood type, randomly assigned, is valid.
It fails whenever the blood type is O+ or O-,
as the LLM assumed these would be named using zero, rather than
the letter O.
By repeating the test a few times, CoverUp is likely to find
the error and give the LLM a chance to correct it.
While CoverUp does not attempt to detect flaky tests by default, it
supports passing custom arguments to pytest.
This approach lets the user specify that each test should be repeated a number
of times using the pytest-repeat plugin.
Flaky tests are then more likely to fail, leading CoverUp to give the
LLM a chance to correct the problem.
III-FOther Technical Challenges
Making CoverUp a practical tool poses several technical challenges.
One such challenge stems from the time spent in LLM inference and executing tests.
Each individual prompt sent through OpenAI’s API typically requires several seconds
to complete.
Similarly, executing individual LLM-generated tests and measuring the new coverage
achieved requires additional time.
CoverUp repeats this process for each code segment as well as
each time the dialogue is continued.
Even though CoverUp limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, CoverUp would
take an unacceptable amount of time creating tests for packages of nontrivial size.
CoverUp instead prompts for tests and verifies them asynchronously, using the Python
asyncio package.
As a result, during CoverUp’s evaluation (Section IV) we observe
a 500x speedup over serial execution.
Other practical challenges arise because OpenAI’s API is
provided as a cloud service and subject to various limits.
CoverUp handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, CoverUp spreads out its
requests using a leaky bucket [29] scheme implemented by the
Python module aiolimiter.
Additionally, CoverUp can create and use checkpoint files, allowing the user
to resume prompting for tests (and not lose any progress so far) after
interrupting it or stopping due to unforeseen circumstances, such as the
OpenAI account running out of funds.
IV Evaluation
Our evaluation investigates the following questions:
RQ1:
How does CoverUp’s coverage compare to the previous state of the art?
(Section IV-B)
RQ2:
How does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
(Section IV-C)
RQ3:
How effective is CoverUp’s continued dialogue at increasing coverage?
(Section IV-D)
IV-AExperimental Setup
Benchmarks
We utilize a benchmark suite collated by the authors
of CodaMosa [2], available at https://github.com/microsoft/codamosa;
we refer to it as suite CM.
It stems originally from 35 projects used in the evaluation of
BugsInPy [30] and
Pynguin [31].
The suite is made up of a subset of the Python modules in these projects, selected
to focus on cases where Pynguin has difficulty achieving full coverage:
after removing the modules that failed to produce results, removing the modules
where Pynguin achieved full coverage within one minute, and downsampling the
number of modules with a common parent, the suite contains 486 benchmark modules
from 27 projects.
Given suite CM’s focus, evaluating CoverUp only on it would leave open
the question of how it performs on code where SBST, or more precisely, the
original pre-CodaMosa Pynguin already performed well.
For that reason, to investigate RQ1 we create an additional suite from the same source,
using only modules for which Pynguin achieved full coverage, naming it PY.
Further, if a module is already in CM, indicating that Pynguin failed at least once
to achieve full coverage, we exclude it from PY.
By evaluating CoverUp on both suites, we obtain a fuller
picture of CoverUp’s performance.
Baselines
We use two versions of CodaMosa, the previous state of the art in
automated case test generation, as baselines for evaluation on the CM suite:
the original version that uses the Codex model (“CodaMosa (codex)”) and
our adapted version that uses the same model as CoverUp (“CodaMosa (gpt4)”).
To investigate RQ1, we utilize the tests CodaMosa (codex) originally generated
using its best performing configuration (codamosa-0.8-uninterp),
available from its “dataset” package at https://github.com/microsoft/codamosa-dataset.
The dataset includes 16 runs of CodaMosa for each of the benchmark modules.
Rather than use Pynguin’s coverage measurements from when the tests were generated,
which differs subtly from SlipCover in how it counts branch coverage, as we
describe below,
we re-execute its tests and measure the resulting coverage (Section IV-A).
For RQ2, we adapt CodaMosa to use the same LLM as CoverUp.
The code-davinci-002 (“codex”) LLM used by the original CodaMosa,
which is no longer available, used code completion prompts, where
the LLM, given a code fragment, responds with a suggestion on how to complete it.
CodaMosa (codex) sent that LLM portions of the code under test to provide
it with context as well as the header of a unit test function for that code.
We create CodaMosa (gpt-4) by modifying it to use OpenAI’s chat API and
to insert instructions requesting a code completion before CodaMosa (codex)’s
original prompt.
Figure 10 shows an example of the resulting prompt.
We configure CodaMosa (gpt4) to use OpenAI’s gpt-4-1106-preview LLM,
setting a temperature of 0, the same employed by CoverUp in this evaluation, but
otherwise leave the configuration identical to that of codamosa-0.8-uninterp.
Figure 9 shows the details.
Figure 9: Configuration for CodaMosa (gpt4):
This configuration selects the same LLM and temperature as
CoverUp, but is otherwise based on CodaMosa’s
best performing configuration, codamosa-0.8-uninterp.
12Figure 10: Example prompt after adapting CodaMosa to a chat LLM:
The figure shows a CodaMosa completion prompt, adapted for use
with a chat LLM.
The prompt prefixes instructions to complete the code (1) before
including the original prompt (2).
The code shown is from the flutils package.
Metric
We utilize the line, branch and combined line and branch coverage as metrics,
computing these over all benchmark modules.
We also compute the combined line and branch coverage on a per-module basis.
It is necessary to include both line and branch coverage as metrics because
in Python, branch coverage does not subsume line coverage:
various situations can lead the Python interpreter to throw exceptions,
and when thrown, these exceptions are not recorded as branches in coverage
information.
We measure all coverage using SlipCover [22], including
those for CodaMosa (codex) generated tests.
We do not use Pynguin’s coverage measurements from when CodaMosa (codex)
generated them because Pynguin measures branch coverage differently from
both SlipCover and Python’s de facto standard coverage tool, coverage.py [32].
Before executing a program, the Python interpreter compiles the source code to bytecode.
Pynguin counts the jump instructions used to implement conditionals as branches and thus
measures clause coverage [33, 22].
Figure 11 shows an example with a two-clause conditional
and the resulting bytecode.
Counting such jump instructions as branches increases the total number of possible
“branches” and can thus lead to both inflated and understated coverage percentual
numbers, depending on how many of those were followed during execution.
Figure 11: Python bytecode for two-clause conditional:
Python compiles the two-clause conditional if A and B into
bytecode that includes two jump instructions.
If a coverage tool counts these as branches,
it is really measuring clause coverage.
Execution Environment
We execute both the CodaMosa tests and CoverUp using the codamosa-docker
Docker image available at https://github.com/microsoft/codamosa, modified only
to install SlipCover and to disable its default entrypoint script, allowing easy
execution of other scripts.
The image is based on Debian 11 and includes Python 3.10.2, with which we run all benchmarks.
Prior to each measurement, our scripts install the benchmark-specific package.txt
requirements file distributed with CodaMosa using pip.
Unfortunately, pip fails to install various of these requirements.
Since these failures were originally ignored for CodaMosa, we ignore them as well,
in an effort to replicate the original conditions as closely as possible.
GPT-generated tests often assume that certain Python modules are available on the system.
Before each CoverUp run, we install the modules mock, pytest-datadir,
pytest-mock, responses and hypothesis.
CoverUp supports automatically installing modules required by tests, but our evaluation
does not use that functionality.
CoverUp options
We configure CoverUp to use OpenAI’s recent gpt-4-1106-preview LLM, setting its “temperature” to zero,
and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section III-A)
at its default of 50 lines, and do not automatically repeat test executions
to look for flaky tests (see Section III-E).
To place CoverUp on the same footing as CodaMosa, we first run it without any
pre-existing test suite.
We then run it two more times, allowing it to build upon the test suite from the
previous runs.
We utilize CoverUp’s disable failing test resolution method that
simply disables any tests that fail during integration check
(see Section III-D), and do not pass any extra
pytest flags (see Section III-E).
IV-BComparison to the previous state of the art
To compare CoverUp to the previous state of the art, we first evaluate it on the CM
suite, using CodaMosa (codex) as the baseline.
Figure 12 shows the line, branch and combined line and
branch coverage obtained by CoverUp and the baseline over the entire benchmark suite;
Figure 13 shows the median of those measurements
on a module-by-module basis.
Additionally, Figure 16 plots the difference between the
combined line and branch coverage obtained by CoverUp and CodaMosa on that suite,
also on a module-by-module basis;
green bars in the plot indicate modules where CoverUp achieved higher coverage.
As the figures show, CoverUp achieves higher line, branch and combined
line and branch coverage than CodaMosa (codex) both measuring over the entire
benchmark code base and on a per-module basis.
Figure 12: [RQ1, RQ2] CoverUp yields higher overall coverage:CoverUp yields higher line, branch and line+branch coverage
over the entire benchmark source than both CodaMosa (codex),
and CodaMosa (gpt4).
Figure 13: [RQ1, RQ2] CoverUp yields higher median per-module coverage:CoverUp yields higher median line, branch and line+branch coverage
on a module-by-module basis than than both CodaMosa (codex) and CodaMosa (gpt4).
Figure 14: [RQ1] CoverUp coverage on the PY suite:CoverUp yields near 100% overall and 100% median coverage
on all metrics, showing that its effectiveness is not limited to
the modules selected for the CM suite.
Figure 15: [RQ3] Contribution of iterative coverage-guided approach:
almost half (42.9%) of the successes were obtained
by iterating and providing updated coverage information, demonstrating its importance.
Figure 16: [RQ1] CoverUp yields higher coverage than the state of the art:
The graph shows the difference between the lines + branches coverage achieved by
CoverUp and CodaMosa (codex) for the various benchmark modules in the CM benchmark suite.
Green bars indicate modules where CoverUp achieved a higher coverage.
We next evaluate the performance of CoverUp on code that the
original Pynguin already performed well.
Figure 14 shows the results.
Given CoverUp’s near 100% overall coverage and 100% median per-module coverage
results, we conclude that CoverUp also performs well on such code.
[RQ1] Summary:CoverUp achieves significantly higher coverage than CodaMosa (codex),
the previous state of the art.
IV-CHow does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
To investigate RQ2, we run CodaMosa (gpt4) on the CM suite and measure the coverage it obtains.
Figures 12 and 13 show the results,
measuring coverage over the entire benchmark suite and on a per-module basis.
As the figures show, CoverUp also outperforms CodaMosa (gpt4), achieving higher coverage
both across the entire suite and a per-module basis.
We also observe that CodaMosa (gpt4)’s performance differs little from that of
CodaMosa (codex), with overall coverage and per-module line and combined line and
branch measurements within 2% of each other.
[RQ2] Summary:CoverUp still outperforms CodaMosa using a state-of-the-art LLM.
IV-DHow effective are CoverUp’s continued dialogue?
To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
identifying each successful test (i.e., which passes and improves coverage) that
was generated immediately upon the first prompt, after continuing the chat with a second prompt,
or after a third prompt.
By default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts.
Figure 15 shows the percentage of successes after
the first, second and third prompts.
As the figure shows, while the success rate decreases with each iteration, almost
half of the successes (49.2%) were obtained through iterative refinement of the prompt,
demonstrating its importance.
[RQ3] Summary: continuing the chat contributes to nearly half of successes,
demonstrating its effectiveness.
V Threats to Validity
Benchmark selection
We utilize the same benchmark suite as CodaMosa to facilitate
a direct comparison to the state of the art.
While so far our experience executing CoverUp to generate tests
for other software has yielded similarly high coverage, selecting
a different set of benchmarks could yield different results.
Execution environment
CodaMosa’s original evaluation environment failed to install various
Python modules that are prerequisites to the applications used as benchmarks.
In an effort to replicate the original conditions as closely as possible,
we ignored these failures as well.
It seems likely that both CodaMosa and CoverUp would be better able
to generate tests if these modules were not missing.
For CoverUp, the modules contributed by the ansible package are
the most affected, with 326 test generation attempts being aborted due to
missing modules.
LLM model dependency
All of CoverUp’s development and evaluation was performed using
OpenAI GPT-4 and GPT-4 Turbo models.
While CoverUp’s test execution and coverage measurement portions
are independent of the LLM, its ultimate performance naturally
depends heavily on the model’s ability to interpret its prompts
and generate tests as requested.
VI Discussion and Future Work
In an attempt to guide the LLM towards useful tests, CoverUp requests
that tests contain assertions in its initial prompt (Section III-B)
and also treats as an error any tests that fail (Section III-C).
It is still possible, however, that the assertions generated by the LLM
are not sufficient to verify the test postconditions.
Even though this paper focuses on case generation and coverage,
rather than the test oracle, we examined a sample of
the assertions in the generated tests.
Many of the assertions seemed appropriate, verifying that the code under test had
the intended effect, but others, for example, merely checked that the
state had changed.
In future work, we intend to extend CoverUp to evaluate the assertions,
possibly employing mutation testing [34, 10],
and contining the dialogue (Section III-C) to request improvements.
Another case that is likely to be a good candidate for continued prompting is
when the LLM generates a test requiring Python modules that are not present
on the system.
As we describe in Section III-C, CoverUp detects this situation
and optionally attempts to install or records a list of missing modules for
later installation; it could, however, instead prompt the LLM to adjust or
re-formulate the test so that the module is not necessary.
Initial experiments suggest that this may be a viable way to automatically
resolve such cases.
A number of initial prompt test generation failures could be attributed
to CoverUp not providing enough context in its prompts, such as related
definitions from other source files.
In future work, we intend to experiment with ways to provide additional
context.
In particular, OpenAI’s function calling feature, which allows
a model to request additional information in response to a prompt, may
work well for this purpose, reducing the need for continued prompting.
Finally, we developed and evaluated CoverUp exclusively using OpenAI
models (GPT-4, GPT-4 Turbo); we intend to adapt CoverUp to work with
other models as well.
We believe that adapting it for use with on-premises LLMs might be of
special interest to industry practitioners, as it would help protect
intellectual property.
VII Conclusion
This paper presents CoverUp, a novel approach for guiding
LLM-based test generation with coverage analysis.
We show that coupling coverage information with prompting, and interatively
refining prompts with updated coverage information to focus the LLM’s attention
on improving uncovered code, is effective at generating test suites with
higher coverage than past approaches.
test case generator
line
branch
line + branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4)
TABLE I: [RQ1, RQ2] Overall coverage obtained on the CM suite.
test case generator
line
branch
line + branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4)
TABLE II: [RQ1, RQ2] Median per-module coverage obtained on the CM suite.
metric
line
branch
line + branch
overall coverage
per-module median
TABLE III: [RQ1] CoverUp coverage on the PY suite.
prompt
contribution
initial
second
third
TABLE IV: [RQ3] Contribution of iterative coverage-guided approach to successes.
[1]
S. Lukasczyk and G. Fraser, “Pynguin: automated unit test generation for
python,” in Proceedings of the ACM/IEEE 44th International Conference
on Software Engineering: Companion Proceedings, ser. ICSE ’22. ACM, May 2022. [Online]. Available:
http://dx.doi.org/10.1145/3510454.3516829
[2]
C. Lemieux, J. P. Inala, S. K. Lahiri, and S. Sen, “Codamosa: Escaping
coverage plateaus in test generation with pre-trained large language
models,” in 2023 IEEE/ACM 45th International Conference on Software
Engineering (ICSE), 2023, pp. 919–931.
[3]
C. Boyapati, S. Khurshid, and D. Marinov, “Korat: automated testing based on
java predicates,” in Proceedings of the 2002 ACM SIGSOFT International
Symposium on Software Testing and Analysis, ser. ISSTA ’02. New York, NY, USA: Association for Computing
Machinery, 2002, p. 123–133. [Online]. Available:
https://doi.org/10.1145/566172.566191
[4]
B. P. Miller, L. Fredriksen, and B. So, “An empirical study of the reliability
of UNIX utilities,” Commun. ACM, vol. 33, no. 12, p. 32–44, dec
1990. [Online]. Available: https://doi.org/10.1145/96267.96279
[5]
A. Fioraldi, A. Mantovani, D. Maier, and D. Balzarotti, “Dissecting American
Fuzzy Lop: A fuzzbench evaluation,” ACM Trans. Softw. Eng.
Methodol., vol. 32, no. 2, mar 2023. [Online]. Available:
https://doi.org/10.1145/3580596
[6]
C. Pacheco, S. K. Lahiri, M. D. Ernst, and T. Ball, “Feedback-directed random
test generation,” in ICSE 2007, Proceedings of the 29th International
Conference on Software Engineering, Minneapolis, MN, USA, May 2007, pp.
75–84.
[7]
P. Godefroid, N. Klarlund, and K. Sen, “Dart: directed automated random
testing,” in Proceedings of the 2005 ACM SIGPLAN Conference on
Programming Language Design and Implementation, ser. PLDI ’05. New York, NY, USA: Association for Computing
Machinery, 2005, p. 213–223. [Online]. Available:
https://doi.org/10.1145/1065010.1065036
[8]
K. Sen, D. Marinov, and G. Agha, “Cute: a concolic unit testing engine for
C,” in Proceedings of the 10th European Software Engineering
Conference Held Jointly with 13th ACM SIGSOFT International Symposium on
Foundations of Software Engineering, ser. ESEC/FSE-13. New York, NY, USA: Association for Computing Machinery,
2005, p. 263–272. [Online]. Available:
https://doi.org/10.1145/1081706.1081750
[10]
G. Fraser and A. Zeller, “Mutation-driven generation of unit tests and
oracles,” IEEE Transactions on Software Engineering, vol. 38, no. 2,
pp. 278–292, 2012.
[11]
G. Fraser and A. Arcuri, “Evosuite: automatic test suite generation for
object-oriented software,” in Proceedings of the 19th ACM SIGSOFT
Symposium and the 13th European Conference on Foundations of Software
Engineering, ser. ESEC/FSE ’11. New
York, NY, USA: Association for Computing Machinery, 2011, p. 416–419.
[Online]. Available: https://doi.org/10.1145/2025113.2025179
[12]
A. Panichella, F. M. Kifetew, and P. Tonella, “Reformulating branch coverage
as a many-objective optimization problem,” in 2015 IEEE 8th
International Conference on Software Testing, Verification and Validation
(ICST), 2015, pp. 1–10.
[13]
——, “Automated test case generation as a many-objective optimisation
problem with dynamic selection of the targets,” IEEE Transactions on
Software Engineering, vol. 44, no. 2, pp. 122–158, 2018.
[14]
M. Tufano, D. Drain, A. Svyatkovskiy, S. K. Deng, and N. Sundaresan, “Unit
test case generation with transformers and focal context,” 2021.
[15]
J. Wang, Y. Huang, C. Chen, Z. Liu, S. Wang, and Q. Wang, “Software testing
with large language models: Survey, landscape, and vision,” 2024.
[16]
M. Schäfer, S. Nadi, A. Eghbali, and F. Tip, “An empirical evaluation of
using large language models for automated unit test generation,” IEEE
Transactions on Software Engineering, vol. 50, no. 1, pp. 85–105, 2024.
[17]
P. Bareiß, B. Souza, M. d’Amorim, and M. Pradel, “Code generation tools
(almost) for free? A study of few-shot, pre-trained language models on
code,” 2022.
[18]
K. Claessen and J. Hughes, “Quickcheck: a lightweight tool for random testing
of haskell programs,” in Proceedings of the Fifth ACM SIGPLAN
International Conference on Functional Programming, ser. ICFP ’00. New York, NY, USA: Association for Computing
Machinery, 2000, p. 268–279. [Online]. Available:
https://doi.org/10.1145/351240.351266
[19]
V. Vikram, C. Lemieux, and R. Padhye, “Can large language models write good
property-based tests?” 2023.
[20]
S. K. Lahiri, S. Fakhoury, A. Naik, G. Sakkas, S. Chakraborty, M. Musuvathi,
P. Choudhury, C. von Veh, J. P. Inala, C. Wang, and J. Gao, “Interactive
code generation via test-driven user-intent formalization,” 2023.
[21]
C. S. Xia, M. Paltenghi, J. Le Tian, M. Pradel, and L. Zhang, “Fuzz4all:
Universal fuzzing with large language models,” Proc. IEEE/ACM ICSE,
2024.
[22]
J. A. Pizzorno and E. D. Berger, “Slipcover: Near Zero-Overhead Code
Coverage for Python,” in Proceedings of ACM SIGSOFT International
Symposium on Software Testing and Analysis (ISSTA 2023), 2023.
[23]
J. White, Q. Fu, S. Hays, M. Sandborn, C. Olea, H. Gilbert, A. Elnashar,
J. Spencer-Smith, and D. C. Schmidt, “A prompt pattern catalog to enhance
prompt engineering with ChatGPT,” 2023.
[24]
A. Gyori, A. Shi, F. Hariri, and D. Marinov, “Reliable testing: detecting
state-polluting tests to prevent test dependency,” in Proceedings of
the 2015 International Symposium on Software Testing and Analysis, ser.
ISSTA 2015. New York, NY, USA:
Association for Computing Machinery, 2015, p. 223–233. [Online]. Available:
https://doi.org/10.1145/2771783.2771793
[25]
G. Fraser and A. Arcuri, “A large-scale evaluation of automated unit test
generation using EvoSuite,” ACM Trans. Softw. Eng. Methodol.,
vol. 24, no. 2, dec 2014. [Online]. Available:
https://doi-org.silk.library.umass.edu/10.1145/2685612
[26]
A. Zeller, “Yesterday, my program worked. Today, it does not. Why?”
SIGSOFT Softw. Eng. Notes, vol. 24, no. 6, p. 253–267, oct 1999.
[Online]. Available: https://doi.org/10.1145/318774.318946
[27]
O. Parry, G. M. Kapfhammer, M. Hilton, and P. McMinn, “A survey of flaky
tests,” ACM Trans. Softw. Eng. Methodol., vol. 31, no. 1, oct 2021.
[Online]. Available: https://doi.org/10.1145/3476105
[28]
W. Lam, S. Winter, A. Wei, T. Xie, D. Marinov, and J. Bell, “A large-scale
longitudinal study of flaky tests,” Proc. ACM Program. Lang., vol. 4,
no. OOPSLA, nov 2020. [Online]. Available:
https://doi.org/10.1145/3428270
[29]
J. Turner, “New directions in communications (or which way to the information
age?),” Comm. Mag., vol. 24, no. 10, p. 8–15, oct 1986. [Online].
Available: https://doi.org/10.1109/MCOM.1986.1092946
[30]
R. Widyasari, S. Q. Sim, C. Lok, H. Qi, J. Phan, Q. Tay, C. Tan, F. Wee, J. E.
Tan, Y. Yieh, B. Goh, F. Thung, H. J. Kang, T. Hoang, D. Lo, and E. L. Ouh,
“BugsInPy: a database of existing bugs in Python programs to enable
controlled testing and debugging studies,” in Proceedings of the 28th
ACM Joint Meeting on European Software Engineering Conference and Symposium
on the Foundations of Software Engineering, ser. ESEC/FSE ’20. ACM, Nov. 2020. [Online]. Available:
http://dx.doi.org/10.1145/3368089.3417943
[31]
S. Lukasczyk, F. Kroiß, and G. Fraser, “An empirical study of automated
unit test generation for python,” CoRR, vol. abs/2111.05003, 2021.
[Online]. Available: https://arxiv.org/abs/2111.05003
[33]
N. Li, X. Meng, J. Offutt, and L. Deng, “Is bytecode instrumentation as good
as source code instrumentation: An empirical study with industrial tools
(experience report),” in IEEE 24th International Symposium on
Software Reliability Engineering, ISSRE 2013, Pasadena, CA, USA, November
4-7, 2013. IEEE Computer Society,
2013, pp. 380–389. [Online]. Available:
https://doi.org/10.1109/ISSRE.2013.6698891
[34]
R. DeMillo, R. Lipton, and F. Sayward, “Hints on test data selection: Help for
the practicing programmer,” Computer, vol. 11, no. 4, pp. 34–41,
1978.

### v2

@@
 CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno
-Abstract
-This paper presents CoverUp, a novel system that drives the
-generation of high-coverage Python regression tests via a combination
-of coverage analysis and large-language models (LLMs). CoverUp
-iteratively improves coverage, interleaving coverage analysis with
-dialogs with the LLM to focus its attention on as yet uncovered
-lines and branches. The resulting test suites significantly improve
-coverage over the current state of the art:
-compared to CodaMosa, a hybrid LLM / search-based software
-testing system, CoverUp substantially improves coverage across the board.
-On a per-module basis, CoverUp achieves median line coverage
-of 81% (vs. 62%), branch coverage of 53% (vs. 35%) and
-line+branch coverage of 78% (vs. 55%).
-We show that CoverUp’s iterative, coverage-guided approach is crucial to its effectiveness, contributing to nearly half of its successes.
-††footnotetext: Work done at the University of Massachusetts Amherst.
-I Introduction
-Test generation tools aim to increase the coverage of the program under test.
-By generating tests that cover a wider range of scenarios and execution paths,
+Abstract.
+Testing is an essential part of software development.
+Test generation tools attempt to automate the otherwise labor-intensive task of test creation,
+but generating high-coverage
+tests remains a challenge.
+This paper proposes CoverUp, a novel approach to driving the
+generation of high-coverage Python regression tests. CoverUp
+iteratively improves test coverage, interleaving coverage analysis with
+dialogs with the LLM that steer it to refine tests so that they increase coverage of
+lines and branches.
+We evaluate our prototype CoverUp implementation
+across a benchmark of challenging
+code derived from open-source Python projects, and show that
+CoverUp substantially improves on the state of the art.
+Compared to CodaMosa, a hybrid search/LLM-based test generator,
+CoverUp achieves a per-module median line+branch
+coverage of 80% (vs. 47%).
+Compared to MuTAP, a mutation/LLM-based test generator,
+CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
+We show that CoverUp’s iterative, coverage-guided approach is
+crucial to its effectiveness, contributing to nearly 40% of its successes.
+Test Generation, Regression Testing, Large Language Models, Code Coverage
+††ccs: Software and its engineering Software testing and debugging††footnotetext: Work done at the University of Massachusetts Amherst.
+1. Introduction
+Testing is essential to ensuring the quality of software, but
+manually crafting tests can be so labor intensive that developers
+choose not write them (Daka and Fraser, 2014).
+Test generation tools attempt to automate this task, generally working
+under the assumption that the source code is correct.
+By generating tests based on it, they enable regression testing,
+with the goal to prevent future bugs as the software is modified.
+As these tools add tests that cover wider range of scenarios and execution paths,
 they can help uncover bugs missed by manually written tests.
-Unfortunately, current approaches often fail to achieve substantial coverage.
-Pynguin exemplifies the search-based software testing (SBST)
-approach [1].
-Starting from randomly created test cases, SBST employs genetic
-algorithms to mutate the tests in an attempt to increase coverage.
-Unfortunately, its search process can get
-stuck. CodaMosa addresses this problem by tracking
-Pynguin’s progress [2]; when it concludes that the search process has
-stalled, it prompts a large language model (LLM) for a test. It then
-uses that test to re-seed the SBST, allowing it to resume progress.
-While CodaMosa greatly improves Pynguin’s performance and constitutes
-the previous state of the art, the resulting code coverage can still remain
-relatively low (e.g., 34% overall branch coverage;
-see Section IV for details).
-This paper proposes CoverUp, a novel approach that drives LLM-based test
-generation with coverage analysis. CoverUp incorporates detailed
+This paper proposes CoverUp, a novel approach to test generation
+aimed aimed at achieving high coverage.
+Our key insight is that large language models (LLMs) are capable of
+simultaneously reasoning about code and coverage information.
+We leverage this insight in the design of CoverUp.
+CoverUp incorporates detailed
 coverage information into prompts customized to the current state
@@
 prompt, resulting in tests that significantly improve coverage.
-To our knowledge, CoverUp is the first test generation system
-to prompt an LLM for tests specifying the lines and branches lacking coverage
-and that iteratively refines that prompt using coverage analysis.
-We evaluate CoverUp using CodaMosa as a baseline, and on the
-same benchmark applications originally used to evaluate it. Using the
-state-of-the-art OpenAI’s GPT-4 Turbo LLM for both, we find that
-CoverUp yields higher coverage in 61% of the benchmarks,
-significantly increasing all coverage
-metrics, whether computed across all code or aggregated over the
-various benchmarks.
-CoverUp increases median line coverage from 62% to 81%,
-median branch coverage from 35% to 53% and
-median line+branch coverage from 55% to 78%.
-Across all code, CoverUp increases
-line coverage from 54% to 61%, branch coverage from 34% to 43%,
-and line+branch coverage from 49% to 57%. We show that CoverUp’s iterative, coverage-guided approach
-is key to its effectiveness, accounting for roughly half of its successes.
+Our empirical evaluation compares CoverUp to two state of the art
+test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid
+search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a
+a mutation/LLM-based approach.
+We show that CoverUp substantially improves upon the state of the art.
+Across a benchmark of challenging code derived from
+open-source Python projects and using OpenAI’s GPT-4o LLM,
+CoverUp increases coverage on every metric vs. CodaMosa, achieving
+per-module median line+branch coverage of 80% (vs. 47%) and
+overall line+branch coverage of 60% (vs. 45%).
+Compared to MuTAP on one of the benchmark suites it supports,
+CoverUp again achieves greater or equal coverage on every metric,
+with an overall line+branch coverage of 90% (vs. 77%).
 This paper makes the following contributions:
-It presents CoverUp, a novel system that drives the
+It presents CoverUp, a novel approach that drives the
 generation of high-coverage Python regression tests via a combination
 of coverage analysis and large-language models;
-It conducts an empirical analysis of CoverUp, showing that it
+It conducts a empirical analysis of our CoverUp prototype, showing that it
 significantly advances the state of the art.
-II Related Work
+2. Related Work
 Automated test generation is a well-established field of research.
 Among the various proposed methods are
-specification-based [3],
-random [4, 5],
-feedback-directed random [6],
-symbolic execution guided random (“concolic”) [7, 8, 9],
-search-based software testing (SBST) [10, 11, 1, 12, 13]
-and transformer based [14] approaches.
+specification-based (Boyapati et al., 2002),
+random (Miller et al., 1990; Fioraldi et al., 2023),
+feedback-directed random (Pacheco et al., 2007),
+symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008),
+search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018)
+and transformer-based approaches (Tufano et al., 2020).
 The success of large language models on various tasks has motivated
 their application to software engineering;
-Wang et al. survey 102 recent papers using LLMs for software testing [15].
+Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
 This section focuses on previous work most closely related to CoverUp.
-TestPilot prompts an LLM to generate JavaScript unit tests based on the
-function under test’s implementation, its documentation and usage snippets [16].
-Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
-to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
-refine the prompt in case the new tests do not improve coverage.
+SBST approaches
+Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST)
+approach. Starting from randomly created test cases, SBST employs genetic
+algorithms to mutate the tests in an attempt to increase coverage.
+Unfortunately, its search process can get stuck as test mutations repeatedly
+lead to the same execution paths.
+CodaMosa (Lemieux et al., 2023) addresses this problem by tracking
+Pynguin’s progress; when it concludes that the search process has
+stalled, it prompts an LLM for a test. It then uses that test to
+re-seed the SBST, allowing it to resume progress.
+We empirically compare CoverUp to CodaMosa in Section 4.2.
+LLM-based Approaches
 Bareiß et al. study the performance of the Codex LLM on
-Java test case generation, among other code generation tasks [17].
+Java test case generation, among other code generation tasks (Bareiß et al., 2022).
 Its prompts contain the signature of the method under test, an example of test generation,
@@
 or lack of coverage.
-As Section IV-D shows, this iterative dialogue is
+As Section 4.4 shows, this iterative dialogue is
 responsibe for almost 50% of CoverUp’s successful test generation.
 Vikram et al. discuss prompting LLMs based on API documentation
-to generate property-based tests for Python [18, 19].
+to generate property-based tests for Python (Claessen and Hughes, 2000; Vikram et al., 2023).
 CoverUp instead bases its prompting on the source code and
@@
 TiCoder prompts LLMs to generate tests based on
-a natural language description of the intended functionality of code [20].
+a natural language description of the intended functionality of code (Lahiri et al., 2022).
 Rather than facilitate regression testing, however, TiCoder
@@
 intent.
-Fuzz4All [21] uses two separate LLMs to fuzz test programs in various programming
+TestPilot prompts an LLM to generate JavaScript unit tests based on the
+function under test’s implementation, its documentation and usage snippets (Schäfer et al., 2024).
+Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
+to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
+continue the chat if the new tests do not improve coverage, nor does it use tool functions
+to provide the LLM with additional context.
+ChatUniTest (Chen et al., 2024) prompts LLMs to generate Java unit tests.
+Its use of LLMs is limited to prompting with code and for repairs when
+a generated test fails compilation.
+Unlike CoverUp, ChatUniTest does not employ coverage measurements
+to indicate what lines or branches lack coverage, nor does it request
+improvements if the generated tests do not improve coverage.
+Concurrent Work Using LLM-based Approaches
+The approaches below were developed concurrently with CoverUp, which was initially
+posted on GitHub on August 7, 2023:
+Fuzz4All (Xia et al., 2024) uses two separate LLMs to fuzz test programs in various programming
 languages.
@@
 suite’s coverage.
-CodaMosa [2], mentioned in the introduction, prompts
-the Codex LLM for a test when its SBST algorithm gets stuck. Like
-CoverUp, CodaMosa selects code to prompt for based on its (lack
-of) coverage, and the prompt contains source code. However, it does
-not use coverage information within the prompt itself, nor does it
-systematically prompt for tests for all code lacking coverage; it only
-prompts the LLM when necessary to help SBST make progress.
-CodaMosa represents the previous state of the art in test case generation
-for Python, and Section IV uses it as the baseline
-for comparison to CoverUp.
-Figure 1: Overview: The execution steps of CoverUp.
+SymPrompt (Ryan et al., 2024) extracts approximate path constraints from Python functions
+through static analysis and prompts an LLM for tests based on those constraints.
+In contrast, CoverUp prompts an LLM directly indicating the lines and branches that
+lack coverage.
+Whereas SymPrompt extracts related type declarations and other code and include
+these statically in its prompts, CoverUp provides a tool function (OpenAI, 2023)
+that looks up arbitrary additional information about types, functions and other symbols,
+thus allowing the LLM to drive this process.
+SymPrompt’s error handling is limited to deleting lines in the response in an attempt
+to correct syntax errors; CoverUp continues the chat in case of errors, asking the
+LLM for a correction.
+CoverUp also continues the chat if the test does not improve coverage.
+SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546)
+previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
+It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately
+a 2x improvement over its baseline prompt (Ryan et al., 2024).
+Without knowing the exact functions employed despite multiple requests for
+code and/or a replication package111Personal communication with SymPrompt authors.,
+we can only provide a coarse comparison:
+as Section 4.2 shows, using GPT-4o CoverUp
+obtains 82% median per-module line coverage and 64% overall line coverage
+on suite CM.
+The code in that suite was also derived from CodaMosa’s evaluation, but contains
+4,116 functions.
+Even though CoverUp is evaluated on an 5x larger suite, the median per-module
+coverage it achieves is also approximately a 2x improvement over an ablated version
+of itself (Section 4.2).
+TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at
+Meta to improve Kotlin unit tests.
+TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the
+class under test and in some cases the existing test class.
+It discards any tests that do not build, do not pass, or do not improve coverage, and
+presents any remaining tests to a developer for approval during code review.
+Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of
+simply discarding them, it continues the chat with the LLM, asking for improvements.
+CoverUp also indicates in its prompts what portions of the code lack coverage, and
+allow the LLM to discover additional context through a tool function.
+MuTAP (Dakhel et al., 2024) prompts an LLM
+for Python unit tests based on mutation testing.
+MuTAP first prompts for a test for a portion of code and, using it,
+performs mutation testing.
+It then prompts for new assertions for each surviving mutant, adding
+these to the test.
+Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors,
+but, unlike CoverUp, not for other kinds of errors, lack of coverage,
+nor does it include coverage information in its prompts.
+Unfortunately, MuTAP currently lacks any ability to provide context for the code
+under test and hardcodes the datasets used for evaluation in the code,
+hampering efforts to evaluate it on other code.
+In contrast, our implementation of CoverUp, like Pynguin and CodaMosa,
+can be applied to any Python package.
+We empirically compare CoverUp to MuTAP in Section 4.2.
+Figure 1. Graphical Overview of the CoverUp Algorithm.
 After measuring the test suite’s coverage, CoverUp prompts the
@@
 or continuing the dialogue if it needs improvements.
+CoverUp provides a tool function with the prompts with which
+the LLM can request additional context.
 A final check helps ensure that the suite works as a whole.
@@
 to refocus the LLM on any segments missed in previous passes.
-III Technique
-CoverUp first measures the code coverage of the existing test suite
-and splits up the source code into segments that require more testing
-(Section III-A).
-It then prompts the LLM for tests for each of these segments, using a chat
-interface
-(Section III-B).
-As the LLM generates tests, CoverUp executes them, again measuring coverage.
-If the test does not compile, fails or lacks coverage, it continues the chat,
-pointing out any problems and requesting improvements
-(Section III-C).
-Finally, having processed all code segments, CoverUp checks the extended
-test suite, looking for any integration problems
-(Section III-D).
-Figure 1 provides a graphical overview of these steps.
-III-ACode Segmentation
-CoverUp’s first step is to measure the code coverage of any pre-existing test suite;
-it uses SlipCover [22], a recently introduced coverage analyzer that has
-near-zero overhead, for that task.
-It then uses the AST of each source file missing coverage to identify code
-segments that can be used for prompting.
-The goal is to be able to provide to the LLM, with each prompt, a code excerpt
+3. Approach
+This section presents a detailed description of CoverUp.
+Figure 1 provides a graphical overview of its algorithm.
+CoverUp first measures the code coverage obtained by the existing test suite and
+uses it to identify code segments (functions and/or classes) that require additional testing
+(Section 3.1).
+It then prompts an LLM for tests for each of these segments,
+combining information from coverage analysis and static code analysis
+(Section 3.2).
+CoverUp execute the LLM generated tests, once again measuring coverage.
+If the tests do not compile, fail to run or increase coverage, CoverUp continues the
+chat with prompts that request improvements and include any error messages.
+(Section 3.3).
+In both initial and continued prompts, CoverUp offers the LLM a tool function with
+which it can request additional contextual information about names in the source code,
+such as function or type definitions
+(Section 3.4).
+Having handled all code segments, CoverUp checks the extended test suite,
+looking for any integration problems
+(Section 3.5).
+The entire process can be repeated an arbitrary number of times to allow the LLM
+to re-focus on segments not covered by previous iterations.
+3.1. Code Segmentation
+CoverUp’s first step is to measure the code coverage of any pre-existing test suite.
+It then uses the AST of each source file missing coverage to identify code segments
+that need additional coverage.
+Code segments typically consist of a single function or method; subject to a size limit,
+however, they may contain an entire class.
+As Figure 2 shows, if a segment contains a method,
+CoverUp also includes a few lines of the original
+class definition to provide the LLM with context, but omits other methods
+and definitions.
+Since we find that retaining comments improves code comprehension, CoverUp
+does not remove these, thus trading a loss in concision for an increase in generation quality.
+The goal is to provide to the LLM, with each prompt, a code excerpt
 that is intelligible, provides enough context, includes the lines or branches
 lacking coverage, and is as short as possible.
-Keeping code segments short is important for several different reasons.
-Older OpenAI LLMs such as GPT-3.5 and GPT-4, as well as on-premises LLMs like
-Meta’s LLaMA have small input windows of up to a few
-thousand tokens, limiting the maximum size of the code segment.
-OpenAI’s most recent GPT-4 Turbo models allow for inputs of up to 128K tokens,
-but this limit applies not only to the initial prompt, but to the total number of tokens
-in the entire sequence of prompts and responses in the case of a continued chat.
-Finally, OpenAI charges at a per-token rate in both prompts and responses,
-so that it literally pays to be succinct.
-Each code segment typically includes an entire function or method and all its contents.
-If a segment contains a method, CoverUp also includes a few lines of the original
-class definition to provide the LLM with context, but omits other methods
-and definitions; see Figure 2 for an example.
-In an effort to enhance code comprehension, CoverUp retains all comments in a
-segment, although removing them could make the segment more concise.
+Keeping code segments short is important primarily because of LLMs’ limits on
+their context (input) window.
+While various current LLMs such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3 and
+Anthropic’s Claude 3 support large context windows, of over 100K tokens,
+this limit applies not only to the initial prompt, but to the total number of tokens
+in the entire sequence of prompts and responses in the case of tool function calls
+or a continued chat.
+Even when prompts fit in these long context windows, shorter prompts remain preferable:
+Rosas et al. (Rosas et al., 2024) show that
+as the size of prompts containing code increases, so does the likelihood of inaccuracies.
+Finally, LLM API use is often charged at a per-token rate, so that it literally
+pays to be succinct.
 Concretely, to identify the code segments in a source file, CoverUp first
@@
 for another, smaller object containing the line.
-This limit is intended to guide CoverUp towards acceptable excerpt sizes,
-facilitating its use with language models with more limited input window sizes.
 Algorithm 1 describes this process in more detail.
-Figure 2: Example of a code segment with a class method:CoverUp created a code segment for a class method,
-omitting over 200 lines of source code originally
+CoverUp also generates import statements that provide context
+for each code segment: we find that without these, LLMs often make incorrect
+assumptions about symbols in the code, leading to errors.
+Rather than extract these verbatim from source code, CoverUp generates imports
+based on static analysis.
+In particular, it converts any relative imports to absolute imports: using relative
+imports in code excerpts leads to these being used in tests outside of package
+scope, where they are not valid.
+We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced
+coverage analyzer that has near-zero overhead, and Python’s ast module.
+Functionidentify_segments(coverage, max_len)
+foreachindo
+lines_in(missing_branches(coverage, file))
+foreachindo
+empty list
+while( is-a Class) andanddo
+append to
+return
+Algorithm 1Algorithm for identifying the code segments lacking coverage, based on a coverage measurement and a tentative maximum segment length.
+Figure 2. CoverUp summarizes method excerpts:CoverUp generates extremely compact code excerpts.
+In this case, 200 lines of source code were originally
 present between listing lines 1 and 2.
 The original code is from the flutils package.
-III-BInitial Prompting
-CoverUp prompts the LLM for tests for every code segment it identifies;
-Figure 3 shows an example of an initial prompt.
-It has the following structure (circled numbers identify sections in the example):
-a statement assigning the LLM the persona [23]
+3.2. Initial Prompting
+CoverUp next prompts the LLM for tests for each code segment it identified in the previous step.
+Figure 3 shows an example of an initial prompt, with circled numbers
+identifying sections.
+It has the following structure:
+a statement assigning the LLM the persona (White et al., 2023)
 of an “expert Python test-driven developer”, intended to help guide it towards
 high quality tests;
-a sentence pointing out the code excerpt (segment), identifying what file
-and module it comes from, and stating what lines or branches do not execute
-when tested.
+a sentence pointing out the code excerpt (segment), identifying what
+file it comes from, stating what lines or branches do not execute when tested.
 The portion specifying the lines and branches missing coverage is compressed
 using line ranges, simplifying the prompt and reducing token usage.
-a request for a pytest test;
-a series of other requests, such as “make sure that the new test is correct”
-and “include assertions” to steer the result towards usable tests;
+a request for a pytest test functions, and an
+encouragement for the LLM to use the provided tool function;
+a series of other requests, such as
+“include assertions” and “avoid state pollution”, to steer the result towards usable tests;
 a request that the response only include the new Python tests
 to facilitate extracting it from the response and to reduce token usage; and
-the code segment with any relevant lines highlighted by prefixing
-them with their line numbers.
-Relevant lines are those that lack coverage or that are
-part of a branch that lacks coverage.
-23456Figure 3: Example of an initial prompt:
+the code segment, prefixed by generated import statements
+and tagging the lines lacking (line or branch) coverage with their numbers.
+[…]
 The initial prompt selects a persona (1),
@@
 asks for tests (3-5),
-and shows the code segment (6).
-Any lines relevant to the request, i.e., used to identify
-missing coverage, are highlighted by prefixing them with their line numbers
-(as in “19:” in part (6)).
-The code shown is from the flutils package.
-Figure 4: Excerpt of test calling to pytest.main:
-We found that certain tests generated by GPT-4
-included top-level, unconditional calls into pytest.main.
-Such calls, executed when pytest loads the test file,
-caused it to restart the search for all tests in the suite,
-slowing it down immensely.
-Part (4) of the prompt now requests omitting such calls.
-We found that tests generated by GPT-4 often included top-level code calling
+and shows the code segment, prefixed by import statements and
+with the line lacking coverage identified by its number.
+The code shown is from the thonny package.
+We find that tests generated by GPT-4 often include top-level code calling
 into pytest.main or into parts of the test itself.
-Figure 4 shows an example.
 While such top-level code can make sense in a standalone test file, Python
@@
 its test discovery, slowing it down to the point that it became unusable.
-For that reason, part 4 of the prompt includes instructions that the test not
-include such calls.
-Prompt part 4 also instructs that the tests “clean up […], so as not
-to affect other tests”.
-It also suggests the use of pytest-mock, a library useful for isolating
-the software under test from other parts of the code, facilitating test cleanup.
-Despite these instructions, we still observe test generations that include side effects;
-Section III-D describes how CoverUp detects and handles
-such tests.
-As Figure 3 shows, in part 6 of the prompt, CoverUp
-prefixes certain lines with their numbers, such as in 20: before line 20.
-It adds such prefixes to all lines lacking coverage as well as to all lines
+For that reason, part 4 of the prompt directs the LLM to
+not include such calls.
+Prompt part 4 also instructs the LLM to generate tests that
+“clean up […], so as to avoid state pollution”.
+It suggests the use of monkeypatch and pytest-mock, useful for
+automatically cleaning up after tests and isolating the software under test from other parts of the code.
+Including these instructions improves the generated tests; nonetheless
+we still observe test generation that includes side effects.
+Section 3.5 describes how CoverUp handles such tests.
+Within the excerpt in part 6, CoverUp tags certain lines with
+their numbers, as in in “19:” in Figure 3.
+It tags all lines lacking coverage as well as all lines
 that are part of a branch that lacks coverage.
-This improves the LLM’s understanding of the missing coverage:
-we found that prompts that only indicated the line number at the beginning of the
-excerpt led to more tests being generated that did not improve coverage.
-If a code segment has no prior coverage, that is, every one of its lines lacks coverage,
-then CoverUp uses a slightly modified initial prompt where part 2 simply
-states “the code does not execute”.
-Typically, the LLM will not generate enough tests to immediately yield high coverage
-in response to this simpler prompt, but this can be improved upon with a subsequent
-execution of CoverUp, where it refocuses the LLM on those lines missed in the first
-iteration.
-III-CVerification and Continued Chat
+This tagging improves the LLM’s understanding of the missing coverage:
+we find that prompts that only indicate the starting line number at the beginning of the
+excerpt lead to more tests being generated that do not improve coverage.
+3.3. Verification and Continued Chat
 Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
-again measuring coverage using SlipCover.
-This step is made practical by SlipCover’s near-zero overhead:
+again measuring coverage.
+In our implementation, this step is made more efficient by SlipCover’s near-zero overhead:
 using coverage.py, the only other alternative tool for Python,
-introduces a median of 180% overhead [22].
+introduces up 260% overhead (Altmayer Pizzorno and Berger, 2023).
 If the new tests pass and increase coverage, CoverUp saves them.
-But, if they do not increase coverage, or if they result in failures or errors,
+If they do not increase coverage, or if they result in failures or errors,
 CoverUp continues the chat session, pointing out the problem(s) and requesting
 improvements.
-Figures 5 and 6 show examples
-of those prompts.
-12Figure 5: Example of a coverage follow-up prompt:CoverUp indicates to the LLM that a line and
+Figures 4 and 5 show examples
+of these prompts.
+[…]
 some branches still weren’t covered (1), asking that it correct the test (2).
-123Figure 6: Example of an error follow-up prompt:CoverUp indicates to the LLM that an error occurred (1),
+[…]
 requests a repair (2) and includes an excerpt of the error messages (3).
@@
 the test that are absent from the system.
-These are typically test helper modules, such as the pytest-ansible plugin
-used to help test the ansible package, but could also be indicative of
-a module required by the program under test that has not yet been installed.
-CoverUp offers options to record the missing modules in a Python requirements.txt
-format to facilitate making them available for a subsequent run.
-III-DIntegration Check
+These missing modules are typically test helper modules, such as the pytest-ansible plugin
+used to help test the ansible package;
+Our implementation offers options to automatically install missing modules and to
+record these in Python’s requirements.txt format to facilitate making
+their use in a subsequent run.
+3.4. Tool Functions
+Tool functions allow an LLM to interact with external tools (OpenAI, 2023).
+CoverUp exposes a get_info tool function that the LLM
+can call to request additional information about any names in the excerpt,
+such as functions, classes, variables or modules.
+It responds to such function calls by performing static analysis and slicing relevant
+portions of the source code.
+Figure 6. Example of a prompt:
+In response to this initial prompt, shown abbreviated, the LLM might
+respond with a get_info("code.A") function call, as requires
+more information on how to instantiate A in tests.
+Figure 7. Example function call response:
+Asked about code.A, CoverUp would perform static analysis and,
+discovering that A is a class, respond with a slice of its definition.
+In this case, it includes its __init__ constructor, which indicates to
+the LLM that A’s constructor takes the initial_x keyword
+argument.
+The slice does not include any of the contents of method m; the LLM could
+subsequently ask for code.A.m to obtain more information on it.
+For example, when given the prompt shown in Figure 6, the LLM
+might ask for information on code.A.
+The function response, shown in Figure 7, indicates that values
+of x must be passed to the class constructor using a keyword argument.
+This response lets the LLM immediately respond with correct tests; we find that if we do not
+include the tool function in this situaiton, the LLM attempts to simply pass values of x to the
+constructor, leading to errors.
+CoverUp makes get_info available both with initial (Section 3.2),
+and continuation prompts (Section 3.3), so that it can help generate
+both better responses and better corrections.
+3.5. Integration Check
 While CoverUp only saves tests that pass when ran by themselves, it is still
 possible that some of these do not clean up properly, or otherwise have
-side effects (“state pollution”) that cause other tests to fail [24].
-This issue is not unique to LLM-based test generation: Fraser et al. [25]
-report side effects on 50% of test classes generated by EvoSuite [11],
+side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
+This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014)
+report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011),
 an SBST test generator.
@@
 symbols used by the code under test.
-To attempt to detect and handle such tests, once CoverUp has processed all
-outstanding code segments, it executes the entire suite in a single session.
-If tests then fail, CoverUp attempts to identify the cause, disabling tests as
-necessary by renaming their files so that pytest ignores them.
-Such tests remain available for review and possible reactivation by the
-user, as they are known to contribute to coverage.
-CoverUp supports two different methods for handling failing tests.
-It can simply disable all failing tests without attempting to ascertain
-the cause of the errors.
-This method is fast and allows the user to move on quickly to prompting
-for more tests, but may lead to a significant loss in coverage as various
-tests are disabled unnecessarily.
-Alternatively, CoverUp can look for the test or tests actually causing the
-failures by successively running subsets of the entire suite, an approach
-adapted from Delta Debugging [26].
-This method requires additional test runs, and thus more time, but is likely
-to reduce the loss in coverage resulting from disabled tests.
-Figure 7: Excerpt of a test with side effects:
+For a test generation approach to be effective, it must thus include some way to
+handle state pollution.
+Our implementation offers three alternatives:
+Execute each test in isolation:
+This approach prevents memory-based state
+pollution by one test from affecting other tests.
+CoverUp implements efficient test isolation based on the Unix fork()
+system call for the pytest framework.
+This option is the default, which Section 4 uses
+for the empirical evaluation of CoverUp.
+Disable polluting tests:
+This approach executes the entire suite and, upon a test failure,
+searches for the polluting tests and disables them.
+CoverUp implements the search by successively reducing and executing subsets
+of the entire suite;
+Disable failing tests:
+This approach may lead to a significant loss in coverage as tests may
+be disabled unnecessarily, and may still leave a polluting test enabled,
+but allows the user to move on quickly to prompting for more tests.
+Any disabled tests remain available for review and possible reactivation by the user.
+Figure 8. Handling state pollution:
 This LLM-generated test overwrites global constants and then,
@@
 other tests, the missing constants cause other tests to fail.
-CoverUp detects these failures during its integration check and
-optionally either disables all failing tests, or searches for the
-test causing the failures.
-III-EHandling Flaky Tests
+Our implementation offers three alternatives to handle these
+situation: it can execute tests in isolation (the default), so that
+and disable the polluting tests, or disable the failing tests.
+3.6. Handling Flaky Tests
 Flaky tests are those that can both pass or fail inconsistently,
-without any changes to the code under test [27, 28].
+without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
 Their unreliable behavior has a variety of causes;
@@
 Consequently, the test fails whenever that blood type is assigned.
-Figure 8: Example of a flaky test:
+Figure 9. Example of a flaky test:
 This test checks that the Person’s
@@
 the letter O.
-By repeating the test a few times, CoverUp is likely to find
+CoverUp executes each test a few times, making it more likely to find
 the error and give the LLM a chance to correct it.
-While CoverUp does not attempt to detect flaky tests by default, it
-supports passing custom arguments to pytest.
-This approach lets the user specify that each test should be repeated a number
-of times using the pytest-repeat plugin.
-Flaky tests are then more likely to fail, leading CoverUp to give the
-LLM a chance to correct the problem.
-III-FOther Technical Challenges
+Our implementation uses the pytest-repeat module to execute each
+newly generated test multiple times, making any flaky tests more likely to fail.
+If a test fails, CoverUp continues the chat, giving the LLM a chance
+to correct the problem.
+3.7. Other Technical Challenges
 Making CoverUp a practical tool poses several technical challenges.
@@
 each time the dialogue is continued.
-Even though CoverUp limits the amount of time spent waiting on responses and
-test executions, if each code segment was processed serially, CoverUp would
-take an unacceptable amount of time creating tests for packages of nontrivial size.
-CoverUp instead prompts for tests and verifies them asynchronously, using the Python
+Even though our implementation limits the amount of time spent waiting on responses and
+test executions, if each code segment was processed serially, it would
+take an unacceptable amount of time to create tests for packages of nontrivial size.
+Instead, it prompts for tests and verifies them asynchronously, using the Python
 asyncio package.
-As a result, during CoverUp’s evaluation (Section IV) we observe
+As a result, during CoverUp’s evaluation (Section 4) we observe
 a 500x speedup over serial execution.
@@
 provided as a cloud service and subject to various limits.
-CoverUp handles a number of timeout and other error conditions automatically.
-To avoid exceeding the rate limits imposed by OpenAI, CoverUp spreads out its
-requests using a leaky bucket [29] scheme implemented by the
+Our implementation handles a number of timeout and other error conditions automatically.
+To avoid exceeding the rate limits imposed by OpenAI, it spreads out its
+requests using a leaky bucket scheme (Turner, 1986) implemented by the
 Python module aiolimiter.
-Additionally, CoverUp can create and use checkpoint files, allowing the user
+Additionally, we implement checkpointing to files, allowing the user
 to resume prompting for tests (and not lose any progress so far) after
@@
 OpenAI account running out of funds.
-IV Evaluation
+4. Evaluation
 Our evaluation investigates the following questions:
-RQ1:
-How does CoverUp’s coverage compare to the previous state of the art?
-(Section IV-B)
-RQ2:
-How does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
-(Section IV-C)
-RQ3:
-How effective is CoverUp’s continued dialogue at increasing coverage?
-(Section IV-D)
-IV-AExperimental Setup
+RQ1::
+Does the coverage of CoverUp’s generated tests improve upon the state of the art?
+(Section 4.2)
+RQ2::
+How effective is CoverUp compared to simply prompting an LLM for tests?
+(Section 4.3)
+RQ3::
+How effective are CoverUp’s continued dialogues at increasing coverage?
+(Section 4.4)
+4.1. Experimental Setup
 Benchmarks
-We utilize a benchmark suite collated by the authors
-of CodaMosa [2], available at https://github.com/microsoft/codamosa;
-we refer to it as suite CM.
-It stems originally from 35 projects used in the evaluation of
-BugsInPy [30] and
-Pynguin [31].
-The suite is made up of a subset of the Python modules in these projects, selected
-to focus on cases where Pynguin has difficulty achieving full coverage:
-after removing the modules that failed to produce results, removing the modules
-where Pynguin achieved full coverage within one minute, and downsampling the
-number of modules with a common parent, the suite contains 486 benchmark modules
-from 27 projects.
-Given suite CM’s focus, evaluating CoverUp only on it would leave open
-the question of how it performs on code where SBST, or more precisely, the
-original pre-CodaMosa Pynguin already performed well.
-For that reason, to investigate RQ1 we create an additional suite from the same source,
-using only modules for which Pynguin achieved full coverage, naming it PY.
-Further, if a module is already in CM, indicating that Pynguin failed at least once
-to achieve full coverage, we exclude it from PY.
-By evaluating CoverUp on both suites, we obtain a fuller
-picture of CoverUp’s performance.
+We utilize three benchmark suites:
+CM, a benchmark suite on which Pynguin struggles to obtain high coverage.
+Originally collated by the authors of CodaMosa (Lemieux et al., 2023) and available
+from https://github.com/microsoft/codamosa, it is derived from 35 open source
+projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and
+Pynguin (Lukasczyk et al., 2023).
+It contains 100,000 lines of code across 425 Python modules.
+PY, a set of modules originally excluded from CodaMosa’s suite because
+Pynguin already performs well on it.
+We evaluate CoverUp on these modules so as not to leave open the question of
+how well it performs on code where Pynguin / SBST already performed well.
+It contains 5,000 lines of code across 84 Python modules.
+MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021),
+originally used in the evaluation of MuTAP (Dakhel et al., 2024).
+We use this dataset to enable comparisons to MuTAP, since its implementation
+lacks the support needed to to run on arbitrary Python packages.
+It contains 2,000 lines of code across 163 functions.
+CM excludes the mimesis, sanic and thef*ck packages
+from the original CodaMosa suite: mimesis is a package for generating random
+data, making it difficult to generate non-flaky tests for it.
+mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa
+generates do not contain any assertions, and thus do not fail in the face of randomly generated values.
+The sanic and thef*ck packages require modules that either are no longer
+available or conflict with CoverUp.
 Baselines
-We use two versions of CodaMosa, the previous state of the art in
-automated case test generation, as baselines for evaluation on the CM suite:
-the original version that uses the Codex model (“CodaMosa (codex)”) and
-our adapted version that uses the same model as CoverUp (“CodaMosa (gpt4)”).
-To investigate RQ1, we utilize the tests CodaMosa (codex) originally generated
-using its best performing configuration (codamosa-0.8-uninterp),
-available from its “dataset” package at https://github.com/microsoft/codamosa-dataset.
-The dataset includes 16 runs of CodaMosa for each of the benchmark modules.
-Rather than use Pynguin’s coverage measurements from when the tests were generated,
-which differs subtly from SlipCover in how it counts branch coverage, as we
-describe below,
-we re-execute its tests and measure the resulting coverage (Section IV-A).
-For RQ2, we adapt CodaMosa to use the same LLM as CoverUp.
-The code-davinci-002 (“codex”) LLM used by the original CodaMosa,
-which is no longer available, used code completion prompts, where
-the LLM, given a code fragment, responds with a suggestion on how to complete it.
-CodaMosa (codex) sent that LLM portions of the code under test to provide
-it with context as well as the header of a unit test function for that code.
-We create CodaMosa (gpt-4) by modifying it to use OpenAI’s chat API and
-to insert instructions requesting a code completion before CodaMosa (codex)’s
-original prompt.
-Figure 10 shows an example of the resulting prompt.
-We configure CodaMosa (gpt4) to use OpenAI’s gpt-4-1106-preview LLM,
-setting a temperature of 0, the same employed by CoverUp in this evaluation, but
-otherwise leave the configuration identical to that of codamosa-0.8-uninterp.
-Figure 9 shows the details.
-Figure 9: Configuration for CodaMosa (gpt4):
-This configuration selects the same LLM and temperature as
-CoverUp, but is otherwise based on CodaMosa’s
-best performing configuration, codamosa-0.8-uninterp.
-12Figure 10: Example prompt after adapting CodaMosa to a chat LLM:
-The figure shows a CodaMosa completion prompt, adapted for use
-with a chat LLM.
-The prompt prefixes instructions to complete the code (1) before
-including the original prompt (2).
-The code shown is from the flutils package.
+To examine RQ1, we compare against two versions of CodaMosa and
+four versions of MuTAP, varying their LLM and prompt type:
+CodaMosa (codex), the original version that uses the Codex LLM.
+Since Codex is no longer available, we use original tests from CodaMosa’s
+evaluation, using its best performing configuration (0.8-uninterp),
+available from https://github.com/microsoft/codamosa-dataset;
+CodaMosa (gpt4o), our adapted version that uses the same model as
+CoverUp, to help rule out differences in performance due to the use of
+different LLMs.
+MuTAP (codex few-shot), the original version that uses the Codex LLM,
+using a “few-shot” prompt, where two examples of test generation are included
+with the prompt.
+Since Codex is no longer available, we use original tests from MuTAP’s
+evaluation, available from https://github.com/ExpertiseModel/MuTAP;
+MuTAP (codex zero-shot), also the original Codex version, but using
+a prompt without examples.
+Here, too, we use original tests from MuTAP’s evaluation;
+MuTAP (gpt4o few-shot), our adapted version that uses the same model
+as CoverUp, with newly generated tests using a “few-shot” prompt;
+MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot”
+prompt.
+To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use
+OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting
+a code completion before its original code completion prompt.
+For MuTAP, we insert a prompt requesting that it respond in the same format as
+a Codex model.
+We were unable to empirically evaluate (Ryan et al., 2024) as at the time of writing
+it remains unavailable: its authors have indicated that it is not yet publicly available.
+For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes
+a nearly identical prompt, except in that it does not specify how the code lacks
+coverage or tags lines lacking coverage, does not include computed import
+statements, does not offer a tool function for additional context, and does not
+continue the chat in case of errors or lack of coverage.
 Metric
@@
 information.
-We measure all coverage using SlipCover [22], including
-those for CodaMosa (codex) generated tests.
-We do not use Pynguin’s coverage measurements from when CodaMosa (codex)
-generated them because Pynguin measures branch coverage differently from
-both SlipCover and Python’s de facto standard coverage tool, coverage.py [32].
-Before executing a program, the Python interpreter compiles the source code to bytecode.
-Pynguin counts the jump instructions used to implement conditionals as branches and thus
-measures clause coverage [33, 22].
-Figure 11 shows an example with a two-clause conditional
-and the resulting bytecode.
-Counting such jump instructions as branches increases the total number of possible
-“branches” and can thus lead to both inflated and understated coverage percentual
-numbers, depending on how many of those were followed during execution.
-Figure 11: Python bytecode for two-clause conditional:
-Python compiles the two-clause conditional if A and B into
-bytecode that includes two jump instructions.
-If a coverage tool counts these as branches,
-it is really measuring clause coverage.
+We run all generated tests in isolation (see Section 3.5)
+and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
 Execution Environment
-We execute both the CodaMosa tests and CoverUp using the codamosa-docker
+We evaluate both CodaMosa and CoverUp using the codamosa-docker
 Docker image available at https://github.com/microsoft/codamosa, modified only
@@
 requirements file distributed with CodaMosa using pip.
-Unfortunately, pip fails to install various of these requirements.
+Unfortunately, pip fails to install a number of these requirements.
 Since these failures were originally ignored for CodaMosa, we ignore them as well,
 in an effort to replicate the original conditions as closely as possible.
-GPT-generated tests often assume that certain Python modules are available on the system.
-Before each CoverUp run, we install the modules mock, pytest-datadir,
-pytest-mock, responses and hypothesis.
-CoverUp supports automatically installing modules required by tests, but our evaluation
-does not use that functionality.
+For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019)
+for mutation testing, which requires older versions of Python.
 CoverUp options
-We configure CoverUp to use OpenAI’s recent gpt-4-1106-preview LLM, setting its “temperature” to zero,
-and do not limit the number of output tokens.
-We leave CoverUp’s target code segment size (see Section III-A)
-at its default of 50 lines, and do not automatically repeat test executions
-to look for flaky tests (see Section III-E).
+We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM,
+setting its “temperature” to zero, and do not limit the number of output tokens.
+We leave CoverUp’s target code segment size (see Section 3.1)
+at its default of 50 lines, and automatically repeat test executions
+to look for flaky tests (see Section 3.6).
 To place CoverUp on the same footing as CodaMosa, we first run it without any
@@
 previous runs.
-We utilize CoverUp’s disable failing test resolution method that
-simply disables any tests that fail during integration check
-(see Section III-D), and do not pass any extra
-pytest flags (see Section III-E).
-IV-BComparison to the previous state of the art
+Since GPT-generated tests often assume the presence of certain Python modules,
+we configure CoverUp to automatically install any such modules if missing.
+4.2. Comparison to the previous state of the art
 To compare CoverUp to the previous state of the art, we first evaluate it on the CM
-suite, using CodaMosa (codex) as the baseline.
-Figure 12 shows the line, branch and combined line and
-branch coverage obtained by CoverUp and the baseline over the entire benchmark suite;
-Figure 13 shows the median of those measurements
-on a module-by-module basis.
+suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
+Figure 10 shows, on the left, the line, branch and combined line and
+branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
+on the right, it shows the median of those measurements on a module-by-module basis.
 Additionally, Figure 12 plots the difference between the
@@
 As the figures show, CoverUp achieves higher line, branch and combined
-line and branch coverage than CodaMosa (codex) both measuring over the entire
+line and branch coverage than both CodaMosa baselines, both measuring over the entire
 benchmark code base and on a per-module basis.
-Figure 12: [RQ1, RQ2] CoverUp yields higher overall coverage:CoverUp yields higher line, branch and line+branch coverage
-over the entire benchmark source than both CodaMosa (codex),
-and CodaMosa (gpt4).
-Figure 13: [RQ1, RQ2] CoverUp yields higher median per-module coverage:CoverUp yields higher median line, branch and line+branch coverage
-on a module-by-module basis than than both CodaMosa (codex) and CodaMosa (gpt4).
-Figure 14: [RQ1] CoverUp coverage on the PY suite:CoverUp yields near 100% overall and 100% median coverage
+Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage,
+49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage;
+on a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage,
+67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
+We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of
+CodaMosa (codex), with overall coverage measurements within 5% of each other.
+As this difference shows, a newer model does not necessarily lead to higher performance.
+We next compare CoverUp to MuTAP, evaluating it on the MT suite.
+Figure 11 shows, on the left, the line, branch and combined line and
+branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
+on the right, it shows the median of those measurements on a module-by-module basis.
+As the figures show, CoverUp achieves higher line, branch and combined line and branch
+coverage than the MuTAP baselines when measuring over the entire benchmark,
+and greater or equal coverage when measuring on a per-module basis.
+Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage,
+89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage;
+on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics,
+except for MuTAP (gtp4o few-shot), which achieves extremely low coverage
+(17%, 0% and 11%).
+We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that
+its original few-shot prompt confuses GPT-4o, leading it to generate tests for
+the examples that make up the “few shot”, rather than for the function under test.
+When this happens, even if the resulting tests run to completion, they do not
+contribute to coverage.
+As the consistently high per-module median coverage values indicate, the MT suite does not
+contain particularly challenging code.
+In fact, the functions in the suite are entirely self contained, only rarely utilize external code,
+and when they do so, from standard libraries, and commonly include type annotations.
+All of these characteristics greatly simplify the task of test generation.
+By contrast, the functions in the CM suite have numerous dependencies, often use external
+code and are generally unannotated.
+Figure 10. [RQ1] CoverUp yields higher overall and median per-module coverage:
+Across the board, CoverUp yields higher coverage than CodaMosa,
+whether measured over the entire suite or on a module-by-module basis.
+Figure 11. [RQ1] CoverUp yields higher overall and median per-module coverage:CoverUp yields higher or equal coverage than MuTAP,
+whether measured over the entire suite or on a module-by-module basis.
+Figure 12. [RQ1] CoverUp yields higher coverage than the state of the art:
+The graph shows the difference between the lines + branches coverage achieved by
+CoverUp and CodaMosa for the various benchmark modules in the CM benchmark suite,
+sorted by percentage of coverage increase.
+Green bars indicate modules where CoverUp achieved a higher coverage.
+Figure 13. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage
 on all metrics, showing that its effectiveness is not limited to
 the modules selected for the CM suite.
-Figure 15: [RQ3] Contribution of iterative coverage-guided approach:
-almost half (42.9%) of the successes were obtained
-by iterating and providing updated coverage information, demonstrating its importance.
-Figure 16: [RQ1] CoverUp yields higher coverage than the state of the art:
-The graph shows the difference between the lines + branches coverage achieved by
-CoverUp and CodaMosa (codex) for the various benchmark modules in the CM benchmark suite.
-Green bars indicate modules where CoverUp achieved a higher coverage.
+Figure 14. [RQ3] Contribution of iterative coverage-guided approach:
+nearly 40% of the successes were obtained
+by iterating and providing coverage and error feedback,
+demonstrating its importance.
 We next evaluate the performance of CoverUp on code that the
@@
 Figure 14 shows the results.
-Given CoverUp’s near 100% overall coverage and 100% median per-module coverage
+Given CoverUp’s 100% median per-module coverage and near 100% overall coverage
 results, we conclude that CoverUp also performs well on such code.
-[RQ1] Summary:CoverUp achieves significantly higher coverage than CodaMosa (codex),
-the previous state of the art.
-IV-CHow does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
-To investigate RQ2, we run CodaMosa (gpt4) on the CM suite and measure the coverage it obtains.
-Figures 12 and 13 show the results,
-measuring coverage over the entire benchmark suite and on a per-module basis.
-As the figures show, CoverUp also outperforms CodaMosa (gpt4), achieving higher coverage
-both across the entire suite and a per-module basis.
-We also observe that CodaMosa (gpt4)’s performance differs little from that of
-CodaMosa (codex), with overall coverage and per-module line and combined line and
-branch measurements within 2% of each other.
-[RQ2] Summary:CoverUp still outperforms CodaMosa using a state-of-the-art LLM.
-IV-DHow effective are CoverUp’s continued dialogue?
+[RQ1] Summary:CoverUp achieves significantly higher coverage than
+both CodaMosa and MuTAP, outperforming the state of the art.
+4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
+Given many LLMs’ near-human performance on various tasks, including software
+engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how
+much CoverUp contributes to performance.
+To address this question, we compare CoverUp to an ablated LLM-only
+version that utilizes a nearly identical prompt, but which lacks all other
+CoverUp functionality (coverage information, precomputed computed
+import statements, tool function for additional context
+and continued chats).
+As Figure 15 shows,
+CoverUp achieves substantially higher line, branch and combined line and branch coverage
+than CoverUp (ablated) both measuring over the entire benchmark code base
+and on a per-module basis.
+Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage,
+49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage;
+on a per-module basis, it achieves 82% (vs. 44%) line coverage,
+67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
+Figure 15. [RQ2] CoverUp contributes significantly to overall performance:
+Across the board, CoverUp yields higher coverage in comparison to
+the ablated (LLM only) version, whether measured over the entire suite
+or on a module-by-module basis.
+[RQ2] Summary:CoverUp contributes significantly to overall performance;
+its performance is not just due to the LLM employed.
+4.4. How effective are CoverUp’s continued dialogues?
 To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
 identifying each successful test (i.e., which passes and improves coverage) that
-was generated immediately upon the first prompt, after continuing the chat with a second prompt,
-or after a third prompt.
-By default, and also in our evaluation, CoverUp continues the conversation for at most
-two additional prompts.
+was generated immediately upon the first prompt or after continuing the chat with a second
+or third prompt
+(by default, and also in our evaluation, CoverUp continues the conversation for at most
+two additional prompts).
 Figure 14 shows the percentage of successes after
 the first, second and third prompts.
-As the figure shows, while the success rate decreases with each iteration, almost
-half of the successes (49.2%) were obtained through iterative refinement of the prompt,
+As the figure shows, while the success rate decreases with each iteration, about
+40% of the successes were obtained through iterative refinement of the prompt,
 demonstrating its importance.
-[RQ3] Summary: continuing the chat contributes to nearly half of successes,
+[RQ3] Summary: continuing the chat contributes almost half its successes,
 demonstrating its effectiveness.
-V Threats to Validity
+5. Threats to Validity
 Benchmark selection
-We utilize the same benchmark suite as CodaMosa to facilitate
-a direct comparison to the state of the art.
+We utilize the CM and PY benchmark suites to evaluate CoverUp’s
+performance on both challenging and less challenging code, as well
+as to facilitate including Codex-based results;
+we utilize the MT benchmark suite because it is supported by MuTAP
+and it, too, facilitates including Codex-based results.
 While so far our experience executing CoverUp to generate tests
@@
 Execution environment
-CodaMosa’s original evaluation environment failed to install various
-Python modules that are prerequisites to the applications used as benchmarks.
+CodaMosa’s original evaluation environment failed to install a number of
+Python modules that are prerequisites for the applications used as benchmarks.
 In an effort to replicate the original conditions as closely as possible,
@@
 to generate tests if these modules were not missing.
-For CoverUp, the modules contributed by the ansible package are
-the most affected, with 326 test generation attempts being aborted due to
-missing modules.
 LLM model dependency
-All of CoverUp’s development and evaluation was performed using
-OpenAI GPT-4 and GPT-4 Turbo models.
+CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo
+and GPT-4o models.
 While CoverUp’s test execution and coverage measurement portions
-are independent of the LLM, its ultimate performance naturally
-depends heavily on the model’s ability to interpret its prompts
-and generate tests as requested.
-VI Discussion and Future Work
-In an attempt to guide the LLM towards useful tests, CoverUp requests
-that tests contain assertions in its initial prompt (Section III-B)
-and also treats as an error any tests that fail (Section III-C).
-It is still possible, however, that the assertions generated by the LLM
-are not sufficient to verify the test postconditions.
-Even though this paper focuses on case generation and coverage,
-rather than the test oracle, we examined a sample of
-the assertions in the generated tests.
-Many of the assertions seemed appropriate, verifying that the code under test had
-the intended effect, but others, for example, merely checked that the
-state had changed.
-In future work, we intend to extend CoverUp to evaluate the assertions,
-possibly employing mutation testing [34, 10],
-and contining the dialogue (Section III-C) to request improvements.
-Another case that is likely to be a good candidate for continued prompting is
-when the LLM generates a test requiring Python modules that are not present
-on the system.
-As we describe in Section III-C, CoverUp detects this situation
-and optionally attempts to install or records a list of missing modules for
-later installation; it could, however, instead prompt the LLM to adjust or
-re-formulate the test so that the module is not necessary.
-Initial experiments suggest that this may be a viable way to automatically
-resolve such cases.
-A number of initial prompt test generation failures could be attributed
-to CoverUp not providing enough context in its prompts, such as related
-definitions from other source files.
-In future work, we intend to experiment with ways to provide additional
-context.
-In particular, OpenAI’s function calling feature, which allows
-a model to request additional information in response to a prompt, may
-work well for this purpose, reducing the need for continued prompting.
-Finally, we developed and evaluated CoverUp exclusively using OpenAI
-models (GPT-4, GPT-4 Turbo); we intend to adapt CoverUp to work with
-other models as well.
-We believe that adapting it for use with on-premises LLMs might be of
-special interest to industry practitioners, as it would help protect
-intellectual property.
-VII Conclusion
+are independent of the LLM
+and, as Section 4.3 shows, CoverUp significantly
+outperforms an ablated version of itself that just relies on the LLM’s capabilities,
+its ultimate performance naturally depends on the model’s ability to
+interpret its prompts and generate tests as requested.
+6. Conclusion
 This paper presents CoverUp, a novel approach for guiding
 LLM-based test generation with coverage analysis.
-We show that coupling coverage information with prompting, and interatively
+We show that coupling coverage information with prompting, and iteratively
 refining prompts with updated coverage information to focus the LLM’s attention
@@
 higher coverage than past approaches.
-test case generator
+Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and
+far outperforms itself when ablated to simply rely on LLM performance.
+overall coverage
+median per-module coverage
+test generator
+line
+branch
+line + branch
 line
@@
 CoverUp
-CodaMosa (codex)
-CodaMosa (gpt4)
-TABLE I: [RQ1, RQ2] Overall coverage obtained on the CM suite.
-test case generator
-line
-branch
-line + branch
-CoverUp
-CodaMosa (codex)
-CodaMosa (gpt4)
-TABLE II: [RQ1, RQ2] Median per-module coverage obtained on the CM suite.
+MuTAP (codex few-shot)
+MuTAP (codex zero-shot)
+MuTAP (gpt4o zero-shot)
+MuTAP (gpt4o few-shot)
+Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
 metric
@@
 branch
-line + branch
-overall coverage
-per-module median
-TABLE III: [RQ1] CoverUp coverage on the PY suite.
+l+b
+module median
+overall
+Table 3. [RQ1] CoverUp’s performance is not limited to “challenging” code: results for PY suite
 prompt
@@
 third
-TABLE IV: [RQ3] Contribution of iterative coverage-guided approach to successes.
-[1]
-S. Lukasczyk and G. Fraser, “Pynguin: automated unit test generation for
-python,” in Proceedings of the ACM/IEEE 44th International Conference
-on Software Engineering: Companion Proceedings, ser. ICSE ’22. ACM, May 2022. [Online]. Available:
-http://dx.doi.org/10.1145/3510454.3516829
-[2]
-C. Lemieux, J. P. Inala, S. K. Lahiri, and S. Sen, “Codamosa: Escaping
-coverage plateaus in test generation with pre-trained large language
-models,” in 2023 IEEE/ACM 45th International Conference on Software
-Engineering (ICSE), 2023, pp. 919–931.
-[3]
-C. Boyapati, S. Khurshid, and D. Marinov, “Korat: automated testing based on
-java predicates,” in Proceedings of the 2002 ACM SIGSOFT International
-Symposium on Software Testing and Analysis, ser. ISSTA ’02. New York, NY, USA: Association for Computing
-Machinery, 2002, p. 123–133. [Online]. Available:
-https://doi.org/10.1145/566172.566191
-[4]
-B. P. Miller, L. Fredriksen, and B. So, “An empirical study of the reliability
-of UNIX utilities,” Commun. ACM, vol. 33, no. 12, p. 32–44, dec
-1990. [Online]. Available: https://doi.org/10.1145/96267.96279
-[5]
-A. Fioraldi, A. Mantovani, D. Maier, and D. Balzarotti, “Dissecting American
-Fuzzy Lop: A fuzzbench evaluation,” ACM Trans. Softw. Eng.
-Methodol., vol. 32, no. 2, mar 2023. [Online]. Available:
-https://doi.org/10.1145/3580596
-[6]
-C. Pacheco, S. K. Lahiri, M. D. Ernst, and T. Ball, “Feedback-directed random
-test generation,” in ICSE 2007, Proceedings of the 29th International
-Conference on Software Engineering, Minneapolis, MN, USA, May 2007, pp.
-75–84.
-[7]
-P. Godefroid, N. Klarlund, and K. Sen, “Dart: directed automated random
-testing,” in Proceedings of the 2005 ACM SIGPLAN Conference on
-Programming Language Design and Implementation, ser. PLDI ’05. New York, NY, USA: Association for Computing
-Machinery, 2005, p. 213–223. [Online]. Available:
+Table 4. [RQ3] Contribution of iterative coverage-guided approach to successes.
+A Large-Scale Evaluation of Automated Unit Test
+Generation Using EvoSuite.
+ACM Trans. Softw. Eng. Methodol.
+24, 2, Article 8
+(December 2014),
+42 pages.
+https://doi.org/10.1145/2685612
+Fraser and Zeller (2012)
+Gordon Fraser and
+Andreas Zeller. 2012.
+Mutation-Driven Generation of Unit Tests and
+Oracles.
+IEEE Transactions on Software Engineering
+38, 2 (2012),
+278–292.
+https://doi.org/10.1109/TSE.2011.93
+Godefroid et al. (2005)
+Patrice Godefroid, Nils
+Klarlund, and Koushik Sen.
+2005.
+DART: directed automated random testing. In
+Proceedings of the 2005 ACM SIGPLAN Conference on
+Programming Language Design and Implementation (Chicago, IL, USA)
+(PLDI ’05). Association for
+Computing Machinery, New York, NY, USA,
+213–223.
 https://doi.org/10.1145/1065010.1065036
-[8]
-K. Sen, D. Marinov, and G. Agha, “Cute: a concolic unit testing engine for
-C,” in Proceedings of the 10th European Software Engineering
-Conference Held Jointly with 13th ACM SIGSOFT International Symposium on
-Foundations of Software Engineering, ser. ESEC/FSE-13. New York, NY, USA: Association for Computing Machinery,
-2005, p. 263–272. [Online]. Available:
+Gruber et al. (2024)
+Martin Gruber,
+Muhammad Firhard Roslan, Owain Parry,
+Fabian Scharnböck, Phil McMinn, and
+Gordon Fraser. 2024.
+Do Automatic Test Generation Tools Generate Flaky
+Tests?. In Proceedings of the IEEE/ACM 46th
+International Conference on Software Engineering (Lisbon, Portugal)
+(ICSE ’24). Association for
+Computing Machinery, New York, NY, USA, Article
+47, 12 pages.
+https://doi.org/10.1145/3597503.3608138
+Gyori et al. (2015)
+Alex Gyori, August Shi,
+Farah Hariri, and Darko Marinov.
+2015.
+Reliable testing: detecting state-polluting tests
+to prevent test dependency. In Proceedings of the
+2015 International Symposium on Software Testing and Analysis (Baltimore,
+MD, USA) (ISSTA 2015).
+Association for Computing Machinery,
+New York, NY, USA, 223–233.
+https://doi.org/10.1145/2771783.2771793
+Hałas (2019)
+Konrad Hałas.
+2019.
+MutPy, a mutation testing tool for
+Python.
+https://github.com/mutpy/mutpyRetrieved on 2024-09-11.
+Lahiri et al. (2022)
+Shuvendu K. Lahiri,
+Aaditya Naik, Georgios Sakkas,
+Piali Choudhury, Curtis von Veh,
+Madanlal Musuvathi, Jeevana Priya Inala,
+Chenglong Wang, and Jianfeng Gao.
+2022.
+Interactive Code Generation via Test-Driven
+User-Intent Formalization.
+CoRR abs/2208.05950
+(2022).
+https://doi.org/10.48550/ARXIV.2208.05950
+arXiv:2208.05950
+Lam et al. (2020)
+Wing Lam, Stefan Winter,
+Anjiang Wei, Tao Xie,
+Darko Marinov, and Jonathan Bell.
+2020.
+A large-scale longitudinal study of flaky tests.
+Proc. ACM Program. Lang.
+4, OOPSLA, Article
+202 (November 2020),
+29 pages.
+https://doi.org/10.1145/3428270
+Lemieux et al. (2023)
+Caroline Lemieux,
+Jeevana Priya Inala, Shuvendu K. Lahiri,
+and Siddhartha Sen. 2023.
+CodaMosa: Escaping Coverage Plateaus in Test
+Generation with Pre-trained Large Language Models. In
+2023 IEEE/ACM 45th International Conference on
+Software Engineering (ICSE). 919–931.
+https://doi.org/10.1109/ICSE48619.2023.00085
+Lukasczyk and Fraser (2022)
+Stephan Lukasczyk and
+Gordon Fraser. 2022.
+Pynguin: Automated Unit Test Generation for
+Python. In 44th IEEE/ACM International
+Conference on Software Engineering: Companion Proceedings, ICSE Companion
+2022, Pittsburgh, PA, USA, May 22-24, 2022.
+ACM/IEEE, 168–172.
+https://doi.org/10.1145/3510454.3516829
+Lukasczyk et al. (2023)
+Stephan Lukasczyk, Florian
+Kroiß, and Gordon Fraser.
+2023.
+An empirical study of automated unit test
+generation for Python.
+Empir. Softw. Eng. 28,
+2 (2023), 36.
+https://doi.org/10.1007/S10664-022-10248-W
+Miller et al. (1990)
+Barton P. Miller, Lars
+Fredriksen, and Bryan So.
+1990.
+An Empirical Study of the Reliability of UNIX
+Utilities.
+Commun. ACM 33,
+12 (1990), 32–44.
+https://doi.org/10.1145/96267.96279
+Pacheco et al. (2007)
+Carlos Pacheco,
+Shuvendu K. Lahiri, Michael D. Ernst,
+and Thomas Ball. 2007.
+Feedback-Directed Random Test Generation. In
+29th International Conference on Software
+Engineering (ICSE 2007), Minneapolis, MN, USA, May 20-26, 2007.
+IEEE Computer Society, 75–84.
+https://doi.org/10.1109/ICSE.2007.37
+Panichella et al. (2015)
+Annibale Panichella,
+Fitsum Meshesha Kifetew, and Paolo
+Tonella. 2015.
+Reformulating Branch Coverage as a Many-Objective
+Optimization Problem. In 2015 IEEE 8th
+International Conference on Software Testing, Verification and Validation
+(ICST). 1–10.
+https://doi.org/10.1109/ICST.2015.7102604
+Panichella et al. (2018)
+Annibale Panichella,
+Fitsum Meshesha Kifetew, and Paolo
+Tonella. 2018.
+Automated Test Case Generation as a Many-Objective
+Optimisation Problem with Dynamic Selection of the Targets.
+IEEE Transactions on Software Engineering
+44, 2 (2018),
+122–158.
+https://doi.org/10.1109/TSE.2017.2663435
+Parry et al. (2021)
+Owain Parry, Gregory M.
+Kapfhammer, Michael Hilton, and Phil
+McMinn. 2021.
+A Survey of Flaky Tests.
+ACM Trans. Softw. Eng. Methodol.
+31, 1, Article 17
+(October 2021), 74 pages.
+https://doi.org/10.1145/3476105
+Rosas et al. (2024)
+Miguel Romero Rosas,
+Miguel Torres Sanchez, and Rudolf
+Eigenmann. 2024.
+Should AI Optimize Your Code? A Comparative Study of
+Current Large Language Models Versus Classical Optimizing Compilers.
+arXiv:2406.12146 [cs.AI]
+https://arxiv.org/abs/2406.12146
+Ryan et al. (2024)
+Gabriel Ryan, Siddhartha
+Jain, Mingyue Shang, Shiqi Wang,
+Xiaofei Ma, Murali Krishna Ramanathan,
+and Baishakhi Ray. 2024.
+Code-Aware Prompting: A Study of Coverage-Guided
+Test Generation in Regression Setting using LLM.
+Proc. ACM Softw. Eng.
+1, FSE (2024),
+951–971.
+https://doi.org/10.1145/3643769
+Schäfer et al. (2024)
+Max Schäfer, Sarah Nadi,
+Aryaz Eghbali, and Frank Tip.
+2024.
+An Empirical Evaluation of Using Large Language
+Models for Automated Unit Test Generation.
+IEEE Transactions on Software Engineering
+50, 1 (2024),
+85–105.
+https://doi.org/10.1109/TSE.2023.3334955
+Sen et al. (2005)
+Koushik Sen, Darko
+Marinov, and Gul Agha. 2005.
+CUTE: a concolic unit testing engine for C. In
+Proceedings of the 10th European Software
+Engineering Conference Held Jointly with 13th ACM SIGSOFT International
+Symposium on Foundations of Software Engineering (Lisbon, Portugal)
+(ESEC/FSE-13). Association for
+Computing Machinery, New York, NY, USA,
+263–272.
 https://doi.org/10.1145/1081706.1081750
-[10]
-G. Fraser and A. Zeller, “Mutation-driven generation of unit tests and
-oracles,” IEEE Transactions on Software Engineering, vol. 38, no. 2,
-pp. 278–292, 2012.
-[11]
-G. Fraser and A. Arcuri, “Evosuite: automatic test suite generation for
-object-oriented software,” in Proceedings of the 19th ACM SIGSOFT
-Symposium and the 13th European Conference on Foundations of Software
-Engineering, ser. ESEC/FSE ’11. New
-York, NY, USA: Association for Computing Machinery, 2011, p. 416–419.
-[Online]. Available: https://doi.org/10.1145/2025113.2025179
-[12]
-A. Panichella, F. M. Kifetew, and P. Tonella, “Reformulating branch coverage
-as a many-objective optimization problem,” in 2015 IEEE 8th
-International Conference on Software Testing, Verification and Validation
-(ICST), 2015, pp. 1–10.
-[13]
-——, “Automated test case generation as a many-objective optimisation
-problem with dynamic selection of the targets,” IEEE Transactions on
-Software Engineering, vol. 44, no. 2, pp. 122–158, 2018.
-[14]
-M. Tufano, D. Drain, A. Svyatkovskiy, S. K. Deng, and N. Sundaresan, “Unit
-test case generation with transformers and focal context,” 2021.
-[15]
-J. Wang, Y. Huang, C. Chen, Z. Liu, S. Wang, and Q. Wang, “Software testing
-with large language models: Survey, landscape, and vision,” 2024.
-[16]
-M. Schäfer, S. Nadi, A. Eghbali, and F. Tip, “An empirical evaluation of
-using large language models for automated unit test generation,” IEEE
-Transactions on Software Engineering, vol. 50, no. 1, pp. 85–105, 2024.
-[17]
-P. Bareiß, B. Souza, M. d’Amorim, and M. Pradel, “Code generation tools
-(almost) for free? A study of few-shot, pre-trained language models on
-code,” 2022.
-[18]
-K. Claessen and J. Hughes, “Quickcheck: a lightweight tool for random testing
-of haskell programs,” in Proceedings of the Fifth ACM SIGPLAN
-International Conference on Functional Programming, ser. ICFP ’00. New York, NY, USA: Association for Computing
-Machinery, 2000, p. 268–279. [Online]. Available:
-https://doi.org/10.1145/351240.351266
-[19]
-V. Vikram, C. Lemieux, and R. Padhye, “Can large language models write good
-property-based tests?” 2023.
-[20]
-S. K. Lahiri, S. Fakhoury, A. Naik, G. Sakkas, S. Chakraborty, M. Musuvathi,
-P. Choudhury, C. von Veh, J. P. Inala, C. Wang, and J. Gao, “Interactive
-code generation via test-driven user-intent formalization,” 2023.
-[21]
-C. S. Xia, M. Paltenghi, J. Le Tian, M. Pradel, and L. Zhang, “Fuzz4all:
-Universal fuzzing with large language models,” Proc. IEEE/ACM ICSE,
+Tillmann and de Halleux (2008)
+Nikolai Tillmann and
+Jonathan de Halleux. 2008.
+Pex-White Box Test Generation for .NET. In
+Tests and Proofs - 2nd International Conference,
+TAP 2008, Prato, Italy, April 9-11, 2008. Proceedings(Lecture Notes in Computer Science,
+Vol. 4966), Bernhard
+Beckert and Reiner Hähnle (Eds.).
+Springer, 134–153.
+https://doi.org/10.1007/978-3-540-79124-9_10
+Tufano et al. (2020)
+Michele Tufano, Dawn
+Drain, Alexey Svyatkovskiy, Shao Kun
+Deng, and Neel Sundaresan.
+2020.
+Unit Test Case Generation with Transformers.
+CoRR abs/2009.05617
+(2020).
+arXiv:2009.05617
+https://arxiv.org/abs/2009.05617
+Turner (1986)
+J. Turner.
+1986.
+New directions in communications (or which way to
+the information age?).
+Comm. Mag. 24,
+10 (October 1986),
+8–15.
+https://doi.org/10.1109/MCOM.1986.1092946
+Vikram et al. (2023)
+Vasudev Vikram, Caroline
+Lemieux, and Rohan Padhye.
+2023.
+Can Large Language Models Write Good Property-Based
+Tests?
+CoRR abs/2307.04346
+(2023).
+https://doi.org/10.48550/ARXIV.2307.04346
+arXiv:2307.04346
+Wang et al. (2024)
+Junjie Wang, Yuchao
+Huang, Chunyang Chen, Zhe Liu,
+Song Wang, and Qing Wang.
 2024.
-[22]
-J. A. Pizzorno and E. D. Berger, “Slipcover: Near Zero-Overhead Code
-Coverage for Python,” in Proceedings of ACM SIGSOFT International
-Symposium on Software Testing and Analysis (ISSTA 2023), 2023.
-[23]
-J. White, Q. Fu, S. Hays, M. Sandborn, C. Olea, H. Gilbert, A. Elnashar,
-J. Spencer-Smith, and D. C. Schmidt, “A prompt pattern catalog to enhance
-prompt engineering with ChatGPT,” 2023.
-[24]
-A. Gyori, A. Shi, F. Hariri, and D. Marinov, “Reliable testing: detecting
-state-polluting tests to prevent test dependency,” in Proceedings of
-the 2015 International Symposium on Software Testing and Analysis, ser.
-ISSTA 2015. New York, NY, USA:
-Association for Computing Machinery, 2015, p. 223–233. [Online]. Available:
-https://doi.org/10.1145/2771783.2771793
-[25]
-G. Fraser and A. Arcuri, “A large-scale evaluation of automated unit test
-generation using EvoSuite,” ACM Trans. Softw. Eng. Methodol.,
-vol. 24, no. 2, dec 2014. [Online]. Available:
-https://doi-org.silk.library.umass.edu/10.1145/2685612
-[26]
-A. Zeller, “Yesterday, my program worked. Today, it does not. Why?”
-SIGSOFT Softw. Eng. Notes, vol. 24, no. 6, p. 253–267, oct 1999.
-[Online]. Available: https://doi.org/10.1145/318774.318946
-[27]
-O. Parry, G. M. Kapfhammer, M. Hilton, and P. McMinn, “A survey of flaky
-tests,” ACM Trans. Softw. Eng. Methodol., vol. 31, no. 1, oct 2021.
-[Online]. Available: https://doi.org/10.1145/3476105
-[28]
-W. Lam, S. Winter, A. Wei, T. Xie, D. Marinov, and J. Bell, “A large-scale
-longitudinal study of flaky tests,” Proc. ACM Program. Lang., vol. 4,
-no. OOPSLA, nov 2020. [Online]. Available:
-https://doi.org/10.1145/3428270
-[29]
-J. Turner, “New directions in communications (or which way to the information
-age?),” Comm. Mag., vol. 24, no. 10, p. 8–15, oct 1986. [Online].
-Available: https://doi.org/10.1109/MCOM.1986.1092946
-[30]
-R. Widyasari, S. Q. Sim, C. Lok, H. Qi, J. Phan, Q. Tay, C. Tan, F. Wee, J. E.
-Tan, Y. Yieh, B. Goh, F. Thung, H. J. Kang, T. Hoang, D. Lo, and E. L. Ouh,
-“BugsInPy: a database of existing bugs in Python programs to enable
-controlled testing and debugging studies,” in Proceedings of the 28th
-ACM Joint Meeting on European Software Engineering Conference and Symposium
-on the Foundations of Software Engineering, ser. ESEC/FSE ’20. ACM, Nov. 2020. [Online]. Available:
-http://dx.doi.org/10.1145/3368089.3417943
-[31]
-S. Lukasczyk, F. Kroiß, and G. Fraser, “An empirical study of automated
-unit test generation for python,” CoRR, vol. abs/2111.05003, 2021.
-[Online]. Available: https://arxiv.org/abs/2111.05003
-[33]
-N. Li, X. Meng, J. Offutt, and L. Deng, “Is bytecode instrumentation as good
-as source code instrumentation: An empirical study with industrial tools
-(experience report),” in IEEE 24th International Symposium on
-Software Reliability Engineering, ISSRE 2013, Pasadena, CA, USA, November
-4-7, 2013. IEEE Computer Society,
-2013, pp. 380–389. [Online]. Available:
-https://doi.org/10.1109/ISSRE.2013.6698891
-[34]
-R. DeMillo, R. Lipton, and F. Sayward, “Hints on test data selection: Help for
-the practicing programmer,” Computer, vol. 11, no. 4, pp. 34–41,
-1978.
+Software Testing With Large Language Models:
+Survey, Landscape, and Vision.
+IEEE Trans. Software Eng.
+50, 4 (2024),
+911–936.
+https://doi.org/10.1109/TSE.2024.3368208
+White et al. (2023)
+Jules White, Quchen Fu,
+Sam Hays, Michael Sandborn,
+Carlos Olea, Henry Gilbert,
+Ashraf Elnashar, Jesse Spencer-Smith,
+and Douglas C. Schmidt. 2023.
+A Prompt Pattern Catalog to Enhance Prompt
+Engineering with ChatGPT.
+arXiv:2302.11382 [cs.SE]
+Widyasari et al. (2020)
+Ratnadira Widyasari,
+Sheng Qin Sim, Camellia Lok,
+Haodi Qi, Jack Phan,
+Qijin Tay, Constance Tan,
+Fiona Wee, Jodie Ethelda Tan,
+Yuheng Yieh, Brian Goh,
+Ferdian Thung, Hong Jin Kang,
+Thong Hoang, David Lo, and
+Eng Lieh Ouh. 2020.
+BugsInPy: a database of existing bugs in Python
+programs to enable controlled testing and debugging studies. In
+ESEC/FSE ’20: 28th ACM Joint European Software
+Engineering Conference and Symposium on the Foundations of Software
+Engineering, Virtual Event, USA, November 8-13, 2020,
+Prem Devanbu, Myra B.
+Cohen, and Thomas Zimmermann (Eds.).
+ACM, 1556–1560.
+https://doi.org/10.1145/3368089.3417943
+Xia et al. (2024)
+Chunqiu Steven Xia, Matteo
+Paltenghi, Jia Le Tian, Michael Pradel,
+and Lingming Zhang. 2024.
+Fuzz4All: Universal Fuzzing with Large Language
+Models. In Proceedings of the IEEE/ACM 46th
+International Conference on Software Engineering (Lisbon, Portugal)
+(ICSE ’24). Association for
+Computing Machinery, New York, NY, USA, Article
+126, 13 pages.
+https://doi.org/10.1145/3597503.3639121

### v3

@@
-CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno
-Abstract.
+
 Testing is an essential part of software development.
-Test generation tools attempt to automate the otherwise labor-intensive task of test creation,
-but generating high-coverage
-tests remains a challenge.
-This paper proposes CoverUp, a novel approach to driving the
-generation of high-coverage Python regression tests. CoverUp
-iteratively improves test coverage, interleaving coverage analysis with
-dialogs with the LLM that steer it to refine tests so that they increase coverage of
-lines and branches.
-We evaluate our prototype CoverUp implementation
-across a benchmark of challenging
-code derived from open-source Python projects, and show that
-CoverUp substantially improves on the state of the art.
-Compared to CodaMosa, a hybrid search/LLM-based test generator,
-CoverUp achieves a per-module median line+branch
-coverage of 80% (vs. 47%).
-Compared to MuTAP, a mutation/LLM-based test generator,
-CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
-We show that CoverUp’s iterative, coverage-guided approach is
-crucial to its effectiveness, contributing to nearly 40% of its successes.
+Test generation tools attempt to automate the otherwise labor-intensive task of test creation, but generating high-coverage tests remains challenging.
+This paper proposes CoverUp, a novel approach to driving the generation of high-coverage Python regression tests.
+CoverUp combines coverage analysis, code context, and feedback in prompts that iteratively guide the LLM to generate tests that improve line and branch coverage.
+We evaluate our prototype CoverUp implementation across a benchmark of challenging code derived from open-source Python projects and show that CoverUp substantially improves on the state of the art.
+Compared to CodaMosa, a hybrid search/LLM-based test generator, CoverUp achieves a per-module median line+branch coverage of 80% (vs. 47%).
+Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
+We also demonstrate that CoverUp’s performance stems not only from the LLM used but from the combined effectiveness of its components.
 Test Generation, Regression Testing, Large Language Models, Code Coverage
-††ccs: Software and its engineering Software testing and debugging††footnotetext: Work done at the University of Massachusetts Amherst.
+††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
 1. Introduction
-Testing is essential to ensuring the quality of software, but
-manually crafting tests can be so labor intensive that developers
-choose not write them (Daka and Fraser, 2014).
-Test generation tools attempt to automate this task, generally working
-under the assumption that the source code is correct.
-By generating tests based on it, they enable regression testing,
-with the goal to prevent future bugs as the software is modified.
-As these tools add tests that cover wider range of scenarios and execution paths,
-they can help uncover bugs missed by manually written tests.
-This paper proposes CoverUp, a novel approach to test generation
-aimed aimed at achieving high coverage.
-Our key insight is that large language models (LLMs) are capable of
-simultaneously reasoning about code and coverage information.
+Testing is essential to ensuring software quality, but manually crafting tests can be so labor-intensive that developers choose not to write them (Daka and Fraser, 2014).
+Test generation tools attempt to automate this task, generally assuming that the source code is correct.
+By generating tests based on it, they enable regression testing, which aims to prevent future bugs as the software is modified.
+As these tools add tests that cover a wider range of scenarios and execution paths, they can help uncover bugs missed by manually written tests.
+This paper proposes CoverUp, a novel approach to test generation aimed at achieving high coverage.
+Our key insight is that large language models (LLMs) can simultaneously reason about code and coverage information.
 We leverage this insight in the design of CoverUp.
-CoverUp incorporates detailed
-coverage information into prompts customized to the current state
-of the test suite, focusing the LLM on code that lacks coverage. Once
-the LLM generates a set of tests, CoverUp executes them and
-measures the resulting coverage. If the LLM-generated tests fail to
-sufficiently improve coverage or fail to run altogether, CoverUp
-continues the dialogue with the LLM, requesting changes to improve
-coverage or fix the error.
-In doing so, CoverUp in effect further refines and clarifies the
-prompt, resulting in tests that significantly improve coverage.
-Our empirical evaluation compares CoverUp to two state of the art
-test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid
-search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a
-a mutation/LLM-based approach.
-We show that CoverUp substantially improves upon the state of the art.
-Across a benchmark of challenging code derived from
-open-source Python projects and using OpenAI’s GPT-4o LLM,
-CoverUp increases coverage on every metric vs. CodaMosa, achieving
-per-module median line+branch coverage of 80% (vs. 47%) and
-overall line+branch coverage of 60% (vs. 45%).
-Compared to MuTAP on one of the benchmark suites it supports,
-CoverUp again achieves greater or equal coverage on every metric,
-with an overall line+branch coverage of 90% (vs. 77%).
+CoverUp incorporates detailed coverage information into prompts customized to the current state of the test suite, focusing the LLM on code that lacks coverage.
+Additionally, it provides a tool function (OpenAI, 2023) that allows the LLM to request additional source code context.
+Once the LLM generates a set of tests, CoverUp executes them and measures the resulting coverage.
+If the LLM-generated tests fail to improve coverage or fail to run altogether, CoverUp continues the dialogue with the LLM, requesting changes to improve coverage or fix the error.
+In doing so, CoverUp, in effect, further refines and clarifies the prompt, resulting in tests that significantly improve coverage.
+Our empirical evaluation compares CoverUp to two state-of-the-art test generation systems: CodaMosa (Lemieux et al., 2023), a hybrid search/LLM-based test generator, and MuTAP (Dakhel et al., 2024), a mutation/LLM-based approach.
+We show that CoverUp substantially improves the state of the art.
+Across a benchmark of challenging code derived from open-source Python projects and using OpenAI’s GPT-4o LLM, CoverUp increases coverage on every metric vs.
+CodaMosa, achieving per-module median line+branch coverage of 80% (vs.
+47%) and overall line+branch coverage of 60% (vs.
+45%).
+Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 90% (vs. 77%).
 This paper makes the following contributions:
-It presents CoverUp, a novel approach that drives the
-generation of high-coverage Python regression tests via a combination
-of coverage analysis and large-language models;
-It conducts a empirical analysis of our CoverUp prototype, showing that it
-significantly advances the state of the art.
+It presents CoverUp, a novel approach that drives the generation of high-coverage Python regression tests via a combination of coverage analysis and large language models;
+It conducts an empirical analysis of our CoverUp prototype, showing that it significantly advances the state of the art;
+It conducts an ablation study, showing that CoverUp’s approach plays a substantial role in the prototype’s performance beyond what can be attributed to the LLM.
 2. Related Work
 Automated test generation is a well-established field of research.
-Among the various proposed methods are
-specification-based (Boyapati et al., 2002),
-random (Miller et al., 1990; Fioraldi et al., 2023),
-feedback-directed random (Pacheco et al., 2007),
-symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008),
-search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018)
-and transformer-based approaches (Tufano et al., 2020).
-The success of large language models on various tasks has motivated
-their application to software engineering;
-Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
+Among the various proposed methods are specification-based (Boyapati et al., 2002), random (Miller et al., 1990; Fioraldi et al., 2023), feedback-directed random (Pacheco et al., 2007), symbolic execution guided random (“concolic”) (Godefroid et al., 2005; Sen et al., 2005; Tillmann and de Halleux, 2008), search-based software testing (SBST) (Fraser and Zeller, 2012; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022; Panichella et al., 2015, 2018) and transformer-based approaches (Tufano et al., 2020).
+The success of large language models on various tasks has motivated their application to software engineering; Wang et al. survey 102 recent papers using LLMs for software testing (Wang et al., 2024).
 This section focuses on previous work most closely related to CoverUp.
 SBST approaches
-Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST)
-approach. Starting from randomly created test cases, SBST employs genetic
-algorithms to mutate the tests in an attempt to increase coverage.
-Unfortunately, its search process can get stuck as test mutations repeatedly
-lead to the same execution paths.
-CodaMosa (Lemieux et al., 2023) addresses this problem by tracking
-Pynguin’s progress; when it concludes that the search process has
-stalled, it prompts an LLM for a test. It then uses that test to
-re-seed the SBST, allowing it to resume progress.
+Pynguin (Lukasczyk and Fraser, 2022) employs a search-based software testing (SBST) approach.
+Starting from randomly created test cases, SBST employs genetic algorithms to mutate the tests, aiming to increase coverage.
+Unfortunately, its search process can get stuck as test mutations repeatedly lead to the same execution paths.
+CodaMosa (Lemieux et al., 2023) addresses this problem by tracking Pynguin’s progress; when it concludes that the search process has stalled, it prompts an LLM for a test.
+It then uses that test to re-seed the SBST, allowing it to resume progress.
 We empirically compare CoverUp to CodaMosa in Section 4.2.
 LLM-based Approaches
-Bareiß et al. study the performance of the Codex LLM on
-Java test case generation, among other code generation tasks (Bareiß et al., 2022).
-Its prompts contain the signature of the method under test, an example of test generation,
-and the body of the method; it discards any tests that do not compile.
-By contrast, CoverUp’s prompts are based on code segments lacking coverage and
-they explicitly request tests to improve it.
-CoverUp also continues the chat with the LLM in case of build failure, failing tests,
-or lack of coverage.
-As Section 4.4 shows, this iterative dialogue is
-responsibe for almost 50% of CoverUp’s successful test generation.
-Vikram et al. discuss prompting LLMs based on API documentation
-to generate property-based tests for Python (Claessen and Hughes, 2000; Vikram et al., 2023).
-CoverUp instead bases its prompting on the source code and
-coverage measurements and, while it accepts property-based tests
-if the LLM generates them, it does not explicitly request them.
-TiCoder prompts LLMs to generate tests based on
-a natural language description of the intended functionality of code (Lahiri et al., 2022).
-Rather than facilitate regression testing, however, TiCoder
-generates tests with the goal of clarifying and formalizing user
-intent.
-TestPilot prompts an LLM to generate JavaScript unit tests based on the
-function under test’s implementation, its documentation and usage snippets (Schäfer et al., 2024).
-Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat
-to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it
-continue the chat if the new tests do not improve coverage, nor does it use tool functions
-to provide the LLM with additional context.
+Bareiß et al. study the performance of the Codex LLM on Java test case generation, among other code generation tasks (Bareiß et al., 2022).
+Its prompts contain the signature of the method under test, an example of test generation, and the method’s body; it discards any tests that do not compile.
+By contrast, CoverUp’s prompts are based on code segments lacking coverage, and they explicitly request tests to improve it.
+CoverUp also continues the chat with the LLM in case of build failure, failing tests, or lack of coverage.
+Section 4.4 shows that this iterative dialogue is responsible for nearly 40% of CoverUp’s successful test generation.
+Vikram et al.
+discuss prompting LLMs based on API documentation to generate property-based tests for Python (Vikram et al., 2023).
+CoverUp instead bases its prompting on the source code and coverage measurements, and while it accepts property-based tests if the LLM generates them, it does not explicitly request them.
+TiCoder prompts LLMs to generate tests based on a natural language description of the intended functionality of code (Lahiri et al., 2022).
+Rather than facilitate regression testing, however, TiCoder generates tests to clarify and formalize user intent.
+TestPilot prompts an LLM to generate JavaScript unit tests based on the function under test’s implementation, its documentation, and usage snippets (Schäfer et al., 2024).
+Like CoverUp, TestPilot checks the tests generated by the LLM and continues the chat to refine the prompt in case of errors, but it does not prompt based on coverage, nor does it continue the chat if the new tests do not improve coverage, nor does it use tool functions to provide the LLM with additional context.
 ChatUniTest (Chen et al., 2024) prompts LLMs to generate Java unit tests.
-Its use of LLMs is limited to prompting with code and for repairs when
-a generated test fails compilation.
-Unlike CoverUp, ChatUniTest does not employ coverage measurements
-to indicate what lines or branches lack coverage, nor does it request
-improvements if the generated tests do not improve coverage.
+Its use of LLMs is limited to prompting with code and for repairs when a generated test fails compilation.
+Unlike CoverUp, ChatUniTest does not employ coverage measurements to indicate what lines or branches lack coverage, nor does it request improvements if the generated tests do not improve coverage.
 Concurrent Work Using LLM-based Approaches
-The approaches below were developed concurrently with CoverUp, which was initially
-posted on GitHub on August 7, 2023:
+The approaches below were developed concurrently with CoverUp, which was initially posted on GitHub on August 7, 2023:
 Fuzz4All (Xia et al., 2024) uses two separate LLMs to fuzz test programs in various programming
 languages.
-The distillation LLM takes in arbitrary user input, such as documentation
-and code examples, and generates prompts for the generation LLM, which
-generates test inputs.
-After the initial user input distillation and prompt selection, Fuzz4All repeatedly
-employs the generation LLM, mutating its prompt to produce additional test inputs.
-While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches
-differ fundamentally: Fuzz4All generates test inputs based on documentation or examples,
-relying on an user-provided test oracle for testing; CoverUp instead works based on
-the source code and coverage measurements, looking to add tests that improve the test
-suite’s coverage.
-SymPrompt (Ryan et al., 2024) extracts approximate path constraints from Python functions
-through static analysis and prompts an LLM for tests based on those constraints.
-In contrast, CoverUp prompts an LLM directly indicating the lines and branches that
-lack coverage.
-Whereas SymPrompt extracts related type declarations and other code and include
-these statically in its prompts, CoverUp provides a tool function (OpenAI, 2023)
-that looks up arbitrary additional information about types, functions and other symbols,
-thus allowing the LLM to drive this process.
-SymPrompt’s error handling is limited to deleting lines in the response in an attempt
-to correct syntax errors; CoverUp continues the chat in case of errors, asking the
-LLM for a correction.
+The distillation LLM takes in arbitrary user input, such as documentation and code examples, and generates prompts for the generation LLM, which produces test inputs.
+After the initial user input distillation and prompt selection, Fuzz4All repeatedly employs the generation LLM, mutating its prompt to produce additional test inputs.
+While Fuzz4All and CoverUp both use LLMs to generate test cases, their approaches differ fundamentally: Fuzz4All generates test inputs based on documentation or examples, relying on a user-provided test oracle for testing; CoverUp instead works based on the source code and coverage measurements, looking to add tests that improve the test suite’s coverage.
+SymPrompt (Ryan et al., 2024) focuses on generating tests that cover hard-to-reach code.
+To do so, it first determines the path constraints needed to reach a certain part of the code and then prompts the LLM for tests based on those constraints.
+CoverUp’s tests can also cover such code, but by prompting based on coverage measurements, a simpler and more direct approach.
+While SymPrompt extracts related declarations and includes them statically in its prompts, CoverUp combines statically generated import declarations with a tool function that enables the LLM to drive the code context discovery process.
+SymPrompt’s error handling is limited to deleting lines in the response in an attempt to correct syntax errors; CoverUp continues the chat in case of errors, asking the LLM for a correction.
 CoverUp also continues the chat if the test does not improve coverage.
-SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546)
-previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
-It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately
-a 2x improvement over its baseline prompt (Ryan et al., 2024).
-Without knowing the exact functions employed despite multiple requests for
-code and/or a replication package111Personal communication with SymPrompt authors.,
-we can only provide a coarse comparison:
-as Section 4.2 shows, using GPT-4o CoverUp
-obtains 82% median per-module line coverage and 64% overall line coverage
-on suite CM.
-The code in that suite was also derived from CodaMosa’s evaluation, but contains
-4,116 functions.
-Even though CoverUp is evaluated on an 5x larger suite, the median per-module
-coverage it achieves is also approximately a 2x improvement over an ablated version
-of itself (Section 4.2).
-TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at
-Meta to improve Kotlin unit tests.
-TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the
-class under test and in some cases the existing test class.
-It discards any tests that do not build, do not pass, or do not improve coverage, and
-presents any remaining tests to a developer for approval during code review.
-Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of
-simply discarding them, it continues the chat with the LLM, asking for improvements.
-CoverUp also indicates in its prompts what portions of the code lack coverage, and
-allow the LLM to discover additional context through a tool function.
-MuTAP (Dakhel et al., 2024) prompts an LLM
-for Python unit tests based on mutation testing.
-MuTAP first prompts for a test for a portion of code and, using it,
-performs mutation testing.
-It then prompts for new assertions for each surviving mutant, adding
-these to the test.
-Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors,
-but, unlike CoverUp, not for other kinds of errors, lack of coverage,
-nor does it include coverage information in its prompts.
-Unfortunately, MuTAP currently lacks any ability to provide context for the code
-under test and hardcodes the datasets used for evaluation in the code,
-hampering efforts to evaluate it on other code.
-In contrast, our implementation of CoverUp, like Pynguin and CodaMosa,
-can be applied to any Python package.
+SymPrompt is evaluated on an undisclosed subset of 897 functions (from at least 4,546) previously used by CodaMosa and BugsInPy (Widyasari et al., 2020).
+It achieves 74% line coverage using OpenAI’s GPT-4 LLM, approximately a 2x improvement over its baseline prompt (Ryan et al., 2024).
+Without knowing the exact functions employed despite multiple requests for code and/or a replication package111Personal communication with SymPrompt authors., we can only provide a coarse comparison: as Section 4.2 shows, using GPT-4o CoverUp obtains 82% median per-module line coverage and 64% overall line coverage on a 4,116-function suite also derived from CodaMosa’s evaluation.
+Even though CoverUp is evaluated on a 5x larger suite, the median per-module coverage it achieves is also approximately a 2x improvement over an ablated version of itself (Section 4.2).
+TestGen-LLM (Alshahwan et al., 2024) is a test generation tool developed and deployed at Meta to improve Kotlin unit tests.
+TestGen-LLM prompts LLMs for tests that improve coverage, including in the prompt the class under test and, in some cases, the existing test class.
+It discards any tests that do not build, do not pass, or do not improve coverage and presents any remaining tests to a developer for approval during code review.
+Like TestGen-LLM, CoverUp rejects tests that fail or lack coverage, but instead of simply discarding them, it continues the chat with the LLM, asking for improvements.
+CoverUp also indicates in its prompts what portions of the code lack coverage and allows the LLM to discover additional context through a tool function.
+MuTAP (Dakhel et al., 2024) prompts an LLM for Python unit tests based on mutation testing.
+MuTAP first prompts for a test for a portion of code and, using it, performs mutation testing.
+It then prompts for new assertions for each surviving mutant, adding these to the test.
+Like CoverUp, MuTAP also prompts the LLM for a repair in case of syntax errors, but unlike CoverUp, it does not prompt it for improvements in case of other errors or lack of coverage, nor does it include coverage information in its prompts.
+Unfortunately, MuTAP currently lacks any ability to provide context for the code under test and hardcodes the datasets used for evaluation in the code, hampering efforts to evaluate it on other code.
+In contrast, our implementation of CoverUp, like Pynguin and CodaMosa, can be applied to any Python package.
 We empirically compare CoverUp to MuTAP in Section 4.2.
 Figure 1. Graphical Overview of the CoverUp Algorithm.
-After measuring the test suite’s coverage, CoverUp prompts the
-LLM so as to focus it on code segments that lack coverage.
-It checks each new test, either accepting it if it increases coverage
-or continuing the dialogue if it needs improvements.
-CoverUp provides a tool function with the prompts with which
-the LLM can request additional context.
+After measuring the test suite’s coverage, CoverUp prompts the LLM, focusing it on code segments that lack coverage.
+It checks each new test, either accepting it if it increases coverage or continuing the dialogue if it needs improvements.
+CoverUp provides a tool function with which the LLM can request additional context.
 A final check helps ensure that the suite works as a whole.
-Given its incremental nature, once done CoverUp can be run again
-to refocus the LLM on any segments missed in previous passes.
+Given its incremental nature, once done, CoverUp can be rerun to refocus the LLM on any segments missed in previous passes.
 3. Approach
@@
 Figure 1 provides a graphical overview of its algorithm.
-CoverUp first measures the code coverage obtained by the existing test suite and
-uses it to identify code segments (functions and/or classes) that require additional testing
-(Section 3.1).
-It then prompts an LLM for tests for each of these segments,
-combining information from coverage analysis and static code analysis
-(Section 3.2).
-CoverUp execute the LLM generated tests, once again measuring coverage.
-If the tests do not compile, fail to run or increase coverage, CoverUp continues the
-chat with prompts that request improvements and include any error messages.
-(Section 3.3).
-In both initial and continued prompts, CoverUp offers the LLM a tool function with
-which it can request additional contextual information about names in the source code,
-such as function or type definitions
-(Section 3.4).
-Having handled all code segments, CoverUp checks the extended test suite,
-looking for any integration problems
-(Section 3.5).
-The entire process can be repeated an arbitrary number of times to allow the LLM
-to re-focus on segments not covered by previous iterations.
+CoverUp first measures the code coverage obtained by the existing test suite and uses it to identify code segments (functions and/or classes) that require additional testing (Section 3.1).
+It then prompts an LLM for tests for each segment, combining information from coverage analysis and static code analysis (Section 3.2).
+CoverUp next executes the LLM-generated tests, once again measuring coverage.
+If the tests do not compile, fail to run, or do not increase coverage, CoverUp continues the chat with prompts that request improvements and include any error messages (Section 3.3).
+In both initial and continued prompts, CoverUp offers the LLM a tool function with which it can request additional contextual information about names in the source code, such as function or type definitions (Section 3.4).
+Having handled all code segments, CoverUp checks the extended test suite, looking for any integration problems (Section 3.5).
+The entire process can be repeated arbitrarily, enabling the LLM to refocus on segments not covered in previous iterations.
 3.1. Code Segmentation
 CoverUp’s first step is to measure the code coverage of any pre-existing test suite.
-It then uses the AST of each source file missing coverage to identify code segments
-that need additional coverage.
-Code segments typically consist of a single function or method; subject to a size limit,
-however, they may contain an entire class.
-As Figure 2 shows, if a segment contains a method,
-CoverUp also includes a few lines of the original
-class definition to provide the LLM with context, but omits other methods
-and definitions.
-Since we find that retaining comments improves code comprehension, CoverUp
-does not remove these, thus trading a loss in concision for an increase in generation quality.
-The goal is to provide to the LLM, with each prompt, a code excerpt
-that is intelligible, provides enough context, includes the lines or branches
-lacking coverage, and is as short as possible.
-Keeping code segments short is important primarily because of LLMs’ limits on
-their context (input) window.
-While various current LLMs such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3 and
-Anthropic’s Claude 3 support large context windows, of over 100K tokens,
-this limit applies not only to the initial prompt, but to the total number of tokens
-in the entire sequence of prompts and responses in the case of tool function calls
-or a continued chat.
-Even when prompts fit in these long context windows, shorter prompts remain preferable:
-Rosas et al. (Rosas et al., 2024) show that
-as the size of prompts containing code increases, so does the likelihood of inaccuracies.
-Finally, LLM API use is often charged at a per-token rate, so that it literally
-pays to be succinct.
-Concretely, to identify the code segments in a source file, CoverUp first
-computes a set of “interesting” lines: these are lines that either lack
-coverage, or that are the source or the destination in branches that lack coverage.
-For each of these lines, it then looks in the AST for a class, function or method
-object containing that line.
-If the object found is a class and it spans more lines than a configurable limit,
-CoverUp adds that class definition as segment context and recursively looks
-for another, smaller object containing the line.
+It then uses the AST of each source file missing coverage to identify code segments that need additional coverage.
+Code segments typically consist of a single function or method; subject to a size limit, however, they may contain an entire class.
+As Figure 2 shows, if a segment contains a method, CoverUp also includes a few lines of the original class definition to provide the LLM with context but omits other methods and definitions.
+We find that retaining comments improves code comprehension; for that reason, CoverUp does not remove them, thus trading a loss in concision for an increase in generation quality.
+The goal is to provide the LLM, with each prompt, a code excerpt that is intelligible, provides enough context, includes the lines or branches lacking coverage, and is as short as possible.
+Keeping code segments short is important primarily because of LLMs’ limits on their context (input) window.
+While various current LLMs, such as OpenAI’s GPT-4 and GPT-4o, Meta’s LLaMA 3, and Anthropic’s Claude 3, support large context windows of over 100K tokens, this limit applies to more than just the initial prompt.
+It applies to the number of tokens in the entire sequence of messages included with each request, which grows with each tool function call, LLM response, and continued chat prompt.
+Even when prompts fit in these long context windows, shorter prompts remain preferable: Rosas et al.
+(Rosas et al., 2024) show that as the size of prompts containing code increases, so does the likelihood of inaccuracies.
+Finally, LLM providers typically charge on a per-token basis, so it literally pays to be succinct.
+Concretely, to identify the code segments in a source file, CoverUp first computes a set of “interesting” lines: these are lines that either lack coverage or are the source or the destination in branches that lack coverage.
+It then looks in the AST for a class, function, or method object containing each line.
+If the object found is a class that spans more lines than a configurable limit, CoverUp adds that class definition as segment context and recursively looks for another, smaller object containing the line.
 Algorithm 1 describes this process in more detail.
-CoverUp also generates import statements that provide context
-for each code segment: we find that without these, LLMs often make incorrect
-assumptions about symbols in the code, leading to errors.
-Rather than extract these verbatim from source code, CoverUp generates imports
-based on static analysis.
-In particular, it converts any relative imports to absolute imports: using relative
-imports in code excerpts leads to these being used in tests outside of package
-scope, where they are not valid.
-We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced
-coverage analyzer that has near-zero overhead, and Python’s ast module.
+We implement this step using SlipCover (Altmayer Pizzorno and Berger, 2023), a recently introduced coverage analyzer with near-zero overhead, and Python’s ast module.
 Functionidentify_segments(coverage, max_len)
@@
 Algorithm 1Algorithm for identifying the code segments lacking coverage, based on a coverage measurement and a tentative maximum segment length.
-Figure 2. CoverUp summarizes method excerpts:CoverUp generates extremely compact code excerpts.
-In this case, 200 lines of source code were originally
-present between listing lines 1 and 2.
+Figure 2. CoverUp summarizes method excerpts:CoverUp generates compact code excerpts.
+In this case, 200 lines of source code were originally present between listing lines 1 and 2.
 The original code is from the flutils package.
@@
 CoverUp next prompts the LLM for tests for each code segment it identified in the previous step.
-Figure 3 shows an example of an initial prompt, with circled numbers
-identifying sections.
+Figure 3 shows an example of an initial prompt, with circled numbers identifying sections.
 It has the following structure:
-a statement assigning the LLM the persona (White et al., 2023)
-of an “expert Python test-driven developer”, intended to help guide it towards
-high quality tests;
-a sentence pointing out the code excerpt (segment), identifying what
-file it comes from, stating what lines or branches do not execute when tested.
-The portion specifying the lines and branches missing coverage is compressed
-using line ranges, simplifying the prompt and reducing token usage.
-a request for a pytest test functions, and an
-encouragement for the LLM to use the provided tool function;
-a series of other requests, such as
-“include assertions” and “avoid state pollution”, to steer the result towards usable tests;
-a request that the response only include the new Python tests
-to facilitate extracting it from the response and to reduce token usage; and
+a statement assigning the LLM the persona (White et al., 2023) of an “expert Python test-driven developer”, intended to help guide it towards high quality tests;
+a sentence pointing out the code excerpt (segment), identifying what file it comes from, and stating what lines or branches do not execute when tested.
+The portion specifying the lines and branches missing coverage is compressed using line ranges, simplifying the prompt and reducing token usage.
+a request for pytest test functions and an encouragement for the LLM to use the provided tool function;
+a series of other requests, such as “include assertions” and “avoid state pollution”, to steer the result towards usable tests;
+a request that the response only include the new Python tests to facilitate its extraction from the response and to reduce token usage; and
 the code segment, prefixed by generated import statements
@@
 The code shown is from the thonny package.
-We find that tests generated by GPT-4 often include top-level code calling
-into pytest.main or into parts of the test itself.
-While such top-level code can make sense in a standalone test file, Python
-executes it as part of the loading process and doing so may greatly
-disrupt pytest’s operation.
-In fact, the calls in some of these early tests caused pytest to restart
-its test discovery, slowing it down to the point that it became unusable.
-For that reason, part 4 of the prompt directs the LLM to
-not include such calls.
-Prompt part 4 also instructs the LLM to generate tests that
-“clean up […], so as to avoid state pollution”.
-It suggests the use of monkeypatch and pytest-mock, useful for
-automatically cleaning up after tests and isolating the software under test from other parts of the code.
-Including these instructions improves the generated tests; nonetheless
-we still observe test generation that includes side effects.
+We find that tests generated by GPT-4 often include top-level code calling into pytest.main or into parts of the test itself.
+While such top-level code can make sense in a standalone test file, Python executes it as part of the loading process, and doing so may significantly disrupt pytest’s operation.
+In fact, in some of these early test generations, such calls caused pytest to restart its test discovery, slowing it down until it became unusable.
+For that reason, part 4 of the prompt directs the LLM not to include such calls.
+Prompt part 4 also instructs the LLM to generate tests that “clean up […], so as to avoid state pollution”.
+It suggests using monkeypatch and pytest-mock, useful for automatically cleaning up after tests and isolating the software under test from other parts of the code.
+Including these instructions improves the generated tests; nonetheless, we still observe test generation that includes side effects.
 Section 3.5 describes how CoverUp handles such tests.
-Within the excerpt in part 6, CoverUp tags certain lines with
-their numbers, as in in “19:” in Figure 3.
-It tags all lines lacking coverage as well as all lines
-that are part of a branch that lacks coverage.
-This tagging improves the LLM’s understanding of the missing coverage:
-we find that prompts that only indicate the starting line number at the beginning of the
-excerpt lead to more tests being generated that do not improve coverage.
+Prior to the code excerpt in part 6, CoverUp adds import statements that provide context for each code segment: we find that without these, LLMs often make incorrect assumptions about symbols in the code, leading to errors.
+Rather than extract these verbatim from source code, CoverUp generates imports based on static analysis.
+In particular, it converts any relative imports to absolute imports.
+Using relative imports in code excerpts can lead to tests that use them outside of package scope, where they are invalid.
+Within the code excerpt, CoverUp tags certain lines with their numbers, as in “19:” in Figure 3.
+It tags all lines lacking coverage and those that are part of a branch lacking coverage.
+This tagging improves the LLM’s understanding of the missing coverage: we find that prompts that only indicate the starting line number at the beginning of the excerpt lead to tests that do not improve coverage.
+To send the prompt using OpenAI’s API, CoverUp embeds it as a “message” in a JSON-formatted request that also includes other fields specifying the model to use, meta-parameters such as the model temperature, etc.
+Figure 4 shows an example.
+Figure 4. Sending a prompt: to prompt and LLM using OpenAI’s chat API, CoverUp puts together a JSON-formatted request that includes the prompt in “messages”.
+[...] indicates a portion omitted for brevity.
 3.3. Verification and Continued Chat
-Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
-again measuring coverage.
-In our implementation, this step is made more efficient by SlipCover’s near-zero overhead:
-using coverage.py, the only other alternative tool for Python,
-introduces up 260% overhead (Altmayer Pizzorno and Berger, 2023).
+Once the LLM generates tests in response to the initial prompt, CoverUp executes them, again measuring coverage.
+In our implementation, this step is made more efficient by SlipCover’s near-zero overhead: using coverage.py, the only other alternative tool for Python, introduces up to 260% overhead (Altmayer Pizzorno and Berger, 2023).
 If the new tests pass and increase coverage, CoverUp saves them.
-If they do not increase coverage, or if they result in failures or errors,
-CoverUp continues the chat session, pointing out the problem(s) and requesting
-improvements.
-Figures 4 and 5 show examples
-of these prompts.
+If, conversely, they do not increase coverage or result in failures or errors, CoverUp continues the chat session, pointing out the problem(s) and requesting improvements.
+To continue the chat, CoverUp sends another request to the LLM, appending the LLM’s response and a new prompt to the previous messages.
+Figure 5 shows how CoverUp continues the chat; Figures 6 and 7 show examples of prompts requesting improvements.
+Figure 5.
+Continuing the chat: to continue a chat, CoverUp sends out a new request, including the previous messages (here, the initial prompt), the LLM’s response (indicated with the “assistant” role), and the new prompt.
+[...] indicates a portion omitted for brevity.
 […]
@@
 […]
-requests a repair (2) and includes an excerpt of the error messages (3).
-In the original execution from which this example is taken,
-the LLM responds with an usable test (not shown).
-Prior to executing a new test, CoverUp looks for any Python modules used by
-the test that are absent from the system.
-These missing modules are typically test helper modules, such as the pytest-ansible plugin
-used to help test the ansible package;
-Our implementation offers options to automatically install missing modules and to
-record these in Python’s requirements.txt format to facilitate making
-their use in a subsequent run.
+In the original execution from which this example is taken, the LLM responds with a usable test (not shown).
+Before executing a new set of tests, CoverUp looks for any Python modules used by the test that are absent from the system.
+These missing modules are typically test helper modules, such as the pytest-ansible plugin used to help test the ansible package.
+Our implementation offers options to install missing modules automatically and record these in Python’s requirements.txt format to facilitate their use in a subsequent run.
 3.4. Tool Functions
 Tool functions allow an LLM to interact with external tools (OpenAI, 2023).
-CoverUp exposes a get_info tool function that the LLM
-can call to request additional information about any names in the excerpt,
-such as functions, classes, variables or modules.
-It responds to such function calls by performing static analysis and slicing relevant
-portions of the source code.
-Figure 6. Example of a prompt:
-In response to this initial prompt, shown abbreviated, the LLM might
-respond with a get_info("code.A") function call, as requires
-more information on how to instantiate A in tests.
+CoverUp exposes a get_info tool function, which allows the LLM to request additional information about any names in the excerpt, such as types or variables.
+In response, CoverUp provides a portion of the source code that shows the definition of the requested object.
+To indicate support for the get_info tool function, CoverUp includes its description in a “tools” JSON element within each chat request.
+If the LLM needs to invoke this function, it may respond with a list of get_info call requests instead of generating text, each specifying the symbol to retrieve.
+CoverUp then continues the chat by sending a new request that appends the LLM’s call requests and the retrieved results from get_info to the previous messages.
+Figure 8. Example of a prompt likely to lead to a tool function call:
+In response to this initial prompt, the LLM would likely choose to respond with a request to call get_info("code.A"), as it requires more information on how to instantiate A in tests.
+[...] indicates a portion omitted for brevity.
 Figure 9. Example function call response:
-Asked about code.A, CoverUp would perform static analysis and,
-discovering that A is a class, respond with a slice of its definition.
-In this case, it includes its __init__ constructor, which indicates to
-the LLM that A’s constructor takes the initial_x keyword
-argument.
-The slice does not include any of the contents of method m; the LLM could
-subsequently ask for code.A.m to obtain more information on it.
-For example, when given the prompt shown in Figure 6, the LLM
-might ask for information on code.A.
-The function response, shown in Figure 7, indicates that values
-of x must be passed to the class constructor using a keyword argument.
-This response lets the LLM immediately respond with correct tests; we find that if we do not
-include the tool function in this situaiton, the LLM attempts to simply pass values of x to the
-constructor, leading to errors.
-CoverUp makes get_info available both with initial (Section 3.2),
-and continuation prompts (Section 3.3), so that it can help generate
-both better responses and better corrections.
+Asked about code.A, CoverUp performs static analysis and, discovering that A is a class, responds with an abbreviated form of its definition.
+In this case, it includes its __init__ constructor, which indicates to the LLM that A’s constructor takes the initial_x keyword argument.
+The response does not include any of the contents of method x; the LLM could subsequently ask for code.A.x to obtain more information on it.
+The ... in the Python excerpt is literal: it is a valid no-op token in Python, and CoverUp uses it to indicate where code has been removed.
+For example, when given the prompt shown in Figure 8, the LLM might ask for information on code.A.
+The function response, shown in Figure 9, indicates that values of x must be passed to the class constructor using a keyword argument.
+This response lets the LLM immediately respond with correct tests; we find that if we do not include the tool function in this situation, the LLM attempts to pass values of x to the constructor, leading to errors.
+To generate its response, get_info first looks for a function, class, module, or variable matching the requested name.
+Supporting variables is particularly important, as constants and aliases in Python are created using variable assignments.
+Beginning with the module where the excerpt originated, get_info may extend its search to other modules by following import statements.
+Once it locates the definition, get_info constructs a response by extracting relevant portions from each module.
+For many constructs, such as assignments, import statements, or function definitions, it includes the entire definition.
+However, for classes and modules, including their full definitions may result in prohibitively long responses.
+Simply omitting elements, as done in the initial prompt, can lead to missing essential information or cause the LLM to assume such elements do not exist.
+For example, if get_info had entirely omitted property x in Figure 9, there would be no indication of how attribute A.x was accessed.
+Instead, get_info replaces the bodies of such elements with ..., a valid “no-op” token in Python, and indicates the omission in the response.
+CoverUp makes get_info available both with initial (Section 3.2) and continuation prompts (Section 3.3) so that it can help generate both better responses and better corrections.
 3.5. Integration Check
-While CoverUp only saves tests that pass when ran by themselves, it is still
-possible that some of these do not clean up properly, or otherwise have
-side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
-This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014)
-report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011),
-an SBST test generator.
-Figure 8 shows an example where a generated test
-not only modifies, but also, in a misguided attempt to clean up, deletes global
-symbols used by the code under test.
-For a test generation approach to be effective, it must thus include some way to
-handle state pollution.
+While CoverUp only saves tests that pass when run in isolation, it is still possible that some may fail to clean up after themselves or introduce side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
+This issue is not unique to LLM-based test generation: Fraser et al.
+(Fraser and Arcuri, 2014) report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011), an SBST test generator.
+Figure 10 shows an example where a generated test not only modifies but also deletes global symbols used by the code under test, in a misguided attempt to clean up.
+For a test generation approach to be effective, it must thus include some way to handle state pollution.
 Our implementation offers three alternatives:
 Execute each test in isolation:
-This approach prevents memory-based state
-pollution by one test from affecting other tests.
-CoverUp implements efficient test isolation based on the Unix fork()
-system call for the pytest framework.
-This option is the default, which Section 4 uses
-for the empirical evaluation of CoverUp.
+This approach prevents memory-based state pollution by one test from affecting other tests.
+CoverUp implements efficient test isolation based on the Unix fork() system call for the pytest framework.
+This option is the default, which Section 4 uses for the empirical evaluation of CoverUp.
 Disable polluting tests:
-This approach executes the entire suite and, upon a test failure,
-searches for the polluting tests and disables them.
-CoverUp implements the search by successively reducing and executing subsets
-of the entire suite;
+This approach executes the entire suite and, upon a test failure, searches for the polluting tests and disables them.
+CoverUp implements the search by successively reducing and executing subsets of the entire suite;
 Disable failing tests:
-This approach may lead to a significant loss in coverage as tests may
-be disabled unnecessarily, and may still leave a polluting test enabled,
-but allows the user to move on quickly to prompting for more tests.
+This approach may lead to a significant loss in coverage as tests may be disabled unnecessarily and may still leave a polluting test enabled, but it allows the user to move on quickly to prompting for more tests.
 Any disabled tests remain available for review and possible reactivation by the user.
 Figure 10. Handling state pollution:
-This LLM-generated test overwrites global constants and then,
-in a misguided attempt to clean up, deletes them.
-The test succeeds if executed by itself, but when executed along with
-other tests, the missing constants cause other tests to fail.
-Our implementation offers three alternatives to handle these
-situation: it can execute tests in isolation (the default), so that
-and disable the polluting tests, or disable the failing tests.
+This LLM-generated test overwrites global constants and then deletes them in a misguided attempt to clean up.
+The test succeeds if executed by itself, but when executed along with other tests, the missing constants cause other tests to fail.
+Our implementation offers three alternatives to handle this situation: it can execute tests in isolation (the default), disable the polluting tests, or disable the failing tests.
+[...] indicates a portion omitted for brevity.
 3.6. Handling Flaky Tests
-Flaky tests are those that can both pass or fail inconsistently,
-without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
-Their unreliable behavior has a variety of causes;
-in the context of LLM generated tests, they can also in part
-result from not providing the LLM sufficient information.
-Figure 9 shows a test for the
-mimesis data generator package.
-The code under test randomly assigns a blood type to its
-Person object; the LLM, not having been provided with
-the list of valid names, assumes that it includes 0-
-and 0+, when the actually code uses O-
-and O+ (using the letter O, not zero).
+Flaky tests are those that can both pass or fail inconsistently without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
+Their unreliable behavior has a variety of causes; in the context of LLM-generated tests, they can also, in part, result from not providing the LLM with sufficient information.
+Figure 11 shows a test for the mimesis data generator package.
+The code under test randomly assigns a blood type to its Person object; the LLM, not having been provided with the list of valid names, assumes that it includes 0- and 0+ whereas code actually uses O- and O+ (using the letter O, not zero).
 Consequently, the test fails whenever that blood type is assigned.
 Figure 11. Example of a flaky test:
-This test checks that the Person’s
-blood type, randomly assigned, is valid.
-It fails whenever the blood type is O+ or O-,
-as the LLM assumed these would be named using zero, rather than
-the letter O.
-CoverUp executes each test a few times, making it more likely to find
-the error and give the LLM a chance to correct it.
-Our implementation uses the pytest-repeat module to execute each
-newly generated test multiple times, making any flaky tests more likely to fail.
-If a test fails, CoverUp continues the chat, giving the LLM a chance
-to correct the problem.
+This test checks that a randomly assigned blood type is valid.
+It fails whenever the blood type is O+ or O-, as the LLM assumed these would be named using zero rather than the letter O.
+CoverUp executes each test a few times, making the error more likely to surface and thus allowing the LLM to attempt to correct it.
+Our implementation uses the pytest-repeat module to execute each newly generated test multiple times, making any flaky tests more likely to fail.
+If a test fails, CoverUp continues the chat, allowing the LLM to attempt to correct the problem.
 3.7. Other Technical Challenges
@@
 One such challenge stems from the time spent in LLM inference and executing tests.
-Each individual prompt sent through OpenAI’s API typically requires several seconds
-to complete.
-Similarly, executing individual LLM-generated tests and measuring the new coverage
-achieved requires additional time.
-CoverUp repeats this process for each code segment as well as
-each time the dialogue is continued.
-Even though our implementation limits the amount of time spent waiting on responses and
-test executions, if each code segment was processed serially, it would
-take an unacceptable amount of time to create tests for packages of nontrivial size.
-Instead, it prompts for tests and verifies them asynchronously, using the Python
-asyncio package.
-As a result, during CoverUp’s evaluation (Section 4) we observe
-a 500x speedup over serial execution.
-Other practical challenges arise because OpenAI’s API is
-provided as a cloud service and subject to various limits.
-Our implementation handles a number of timeout and other error conditions automatically.
-To avoid exceeding the rate limits imposed by OpenAI, it spreads out its
-requests using a leaky bucket scheme (Turner, 1986) implemented by the
-Python module aiolimiter.
-Additionally, we implement checkpointing to files, allowing the user
-to resume prompting for tests (and not lose any progress so far) after
-interrupting it or stopping due to unforeseen circumstances, such as the
-OpenAI account running out of funds.
+Each individual prompt sent through OpenAI’s API typically requires several seconds to complete.
+Similarly, executing individual LLM-generated tests and measuring the new coverage achieved requires additional time.
+CoverUp repeats this process for each code segment and each time the dialogue is continued.
+Even though our implementation limits the time spent waiting on responses and test executions, if each code segment were processed serially, creating tests for packages of nontrivial size would take an unacceptable amount of time.
+Instead, it prompts for tests and verifies them asynchronously, using the Python asyncio package.
+As a result, during CoverUp’s evaluation (Section 4), we observe a 500x speedup over serial execution.
+Other practical challenges arise because OpenAI’s API is provided as a cloud service and is subject to various limits.
+Our implementation handles various timeout and other error conditions automatically.
+To avoid exceeding the rate limits imposed by OpenAI, it spreads out its requests using a leaky bucket scheme (Turner, 1986) implemented by the Python module aiolimiter.
+Additionally, we implement checkpointing to files, allowing the user to resume prompting for tests (and not lose any progress so far) after interrupting it or stopping due to unforeseen circumstances, such as the OpenAI account running out of funds.
 4. Evaluation
@@
 (Section 4.4)
+RQ4::
+How does the cost of running CoverUp compare to CodaMosa?
+(Section 4.5)
+RQ5::
+How important are CoverUp’s components to its performance?
+(Section 4.6)
 4.1. Experimental Setup
@@
 CM, a benchmark suite on which Pynguin struggles to obtain high coverage.
-Originally collated by the authors of CodaMosa (Lemieux et al., 2023) and available
-from https://github.com/microsoft/codamosa, it is derived from 35 open source
-projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and
-Pynguin (Lukasczyk et al., 2023).
+Collated originally by the authors of CodaMosa (Lemieux et al., 2023) and available from https://github.com/microsoft/codamosa, it is derived from 35 open-source projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and Pynguin (Lukasczyk et al., 2023).
 It contains 100,000 lines of code across 425 Python modules.
-PY, a set of modules originally excluded from CodaMosa’s suite because
-Pynguin already performs well on it.
-We evaluate CoverUp on these modules so as not to leave open the question of
-how well it performs on code where Pynguin / SBST already performed well.
+PY, a set of modules originally excluded from CodaMosa’s suite because Pynguin already performs well on it.
+We evaluate CoverUp on these modules so as not to leave open the question of how well it performs on code where Pynguin / SBST already performed well.
 It contains 5,000 lines of code across 84 Python modules.
-MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021),
-originally used in the evaluation of MuTAP (Dakhel et al., 2024).
-We use this dataset to enable comparisons to MuTAP, since its implementation
-lacks the support needed to to run on arbitrary Python packages.
+MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021), originally used in the evaluation of MuTAP (Dakhel et al., 2024).
+We use this dataset to enable comparisons to MuTAP since its implementation lacks the support needed to run on arbitrary Python packages.
 It contains 2,000 lines of code across 163 functions.
-CM excludes the mimesis, sanic and thef*ck packages
-from the original CodaMosa suite: mimesis is a package for generating random
-data, making it difficult to generate non-flaky tests for it.
-mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa
-generates do not contain any assertions, and thus do not fail in the face of randomly generated values.
-The sanic and thef*ck packages require modules that either are no longer
-available or conflict with CoverUp.
+CM excludes the mimesis, sanic, and thef*ck packages from the original CodaMosa suite: mimesis is a package for generating random data, making it challenging to generate non-flaky tests for it.
+mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa generates do not contain any assertions and thus do not fail in the face of randomly generated values.
+The sanic and thef*ck packages require modules that either are no longer available or conflict with CoverUp.
 Baselines
-To examine RQ1, we compare against two versions of CodaMosa and
-four versions of MuTAP, varying their LLM and prompt type:
+To examine RQ1, we compare against two versions of CodaMosa and four versions of MuTAP, varying their LLM and prompt type:
 CodaMosa (codex), the original version that uses the Codex LLM.
-Since Codex is no longer available, we use original tests from CodaMosa’s
-evaluation, using its best performing configuration (0.8-uninterp),
-available from https://github.com/microsoft/codamosa-dataset;
-CodaMosa (gpt4o), our adapted version that uses the same model as
-CoverUp, to help rule out differences in performance due to the use of
-different LLMs.
-MuTAP (codex few-shot), the original version that uses the Codex LLM,
-using a “few-shot” prompt, where two examples of test generation are included
-with the prompt.
-Since Codex is no longer available, we use original tests from MuTAP’s
-evaluation, available from https://github.com/ExpertiseModel/MuTAP;
+Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp), available from https://github.com/microsoft/codamosa-dataset;
+CodaMosa (gpt4o), our adapted version that uses the same model as CoverUp, to help rule out differences in performance due to the use of different LLMs.
+MuTAP (codex few-shot), the original version that uses the Codex LLM, using a “few-shot” prompt, where two examples of test generation are included with the prompt.
+Since Codex is no longer available, we use original tests from MuTAP’s evaluation, available from https://github.com/ExpertiseModel/MuTAP;
 MuTAP (codex zero-shot), also the original Codex version, but using
@@
 Here, too, we use original tests from MuTAP’s evaluation;
-MuTAP (gpt4o few-shot), our adapted version that uses the same model
-as CoverUp, with newly generated tests using a “few-shot” prompt;
-MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot”
-prompt.
-To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use
-OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting
-a code completion before its original code completion prompt.
-For MuTAP, we insert a prompt requesting that it respond in the same format as
-a Codex model.
-We were unable to empirically evaluate (Ryan et al., 2024) as at the time of writing
-it remains unavailable: its authors have indicated that it is not yet publicly available.
-For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes
-a nearly identical prompt, except in that it does not specify how the code lacks
-coverage or tags lines lacking coverage, does not include computed import
-statements, does not offer a tool function for additional context, and does not
-continue the chat in case of errors or lack of coverage.
+MuTAP (gpt4o few-shot), our adapted version that uses the same model as CoverUp, with newly generated tests using a “few-shot” prompt;
+MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot” prompt.
+To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting a code completion before its original code completion prompt.
+For MuTAP, we insert a prompt requesting that it respond in the same format as a Codex model.
+We could not empirically evaluate (Ryan et al., 2024) as, at the time of writing, it remains unavailable: its authors have indicated that it is not publicly available.
+For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes a nearly identical prompt, except in that it does not specify how the code lacks coverage or tags lines lacking coverage, does not add computed import statements, does not offer a tool function for additional context, and does not continue the chat in case of errors or lack of coverage.
 Metric
-We utilize the line, branch and combined line and branch coverage as metrics,
-computing these over all benchmark modules.
+We utilize the line, branch, and combined line and branch coverage as metrics, computing these for all benchmark modules.
 We also compute the combined line and branch coverage on a per-module basis.
-It is necessary to include both line and branch coverage as metrics because
-in Python, branch coverage does not subsume line coverage:
-various situations can lead the Python interpreter to throw exceptions,
-and when thrown, these exceptions are not recorded as branches in coverage
-information.
-We run all generated tests in isolation (see Section 3.5)
-and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
+It is necessary to include both line and branch coverage as metrics because in Python, branch coverage does not subsume line coverage: various situations can lead the Python interpreter to throw exceptions, and when thrown, these exceptions are not recorded as branches in coverage information.
+We run all generated tests in isolation (see Section 3.5) and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
 Execution Environment
-We evaluate both CodaMosa and CoverUp using the codamosa-docker
-Docker image available at https://github.com/microsoft/codamosa, modified only
-to install SlipCover and to disable its default entrypoint script, allowing easy
-execution of other scripts.
+We evaluate both CodaMosa and CoverUp using the codamosa-docker Docker image available at https://github.com/microsoft/codamosa, modified only to install SlipCover and to disable its default entrypoint script, allowing easy execution of other scripts.
 The image is based on Debian 11 and includes Python 3.10.2, with which we run all benchmarks.
-Prior to each measurement, our scripts install the benchmark-specific package.txt
-requirements file distributed with CodaMosa using pip.
-Unfortunately, pip fails to install a number of these requirements.
-Since these failures were originally ignored for CodaMosa, we ignore them as well,
-in an effort to replicate the original conditions as closely as possible.
-For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019)
-for mutation testing, which requires older versions of Python.
+As the host system for Docker, we utilize a Linux kernel 5.6 system with 10 Intel i9 cores at 3.7GHz and 64GB RAM.
+Before each measurement, our scripts install the benchmark-specific package.txt requirements file distributed with CodaMosa using pip.
+Unfortunately, pip fails to install some of these requirements.
+Since these failures were originally ignored for CodaMosa, we also ignore them to replicate the original conditions as closely as possible.
+For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019) for mutation testing, which requires older versions of Python.
 CoverUp options
-We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM,
-setting its “temperature” to zero, and do not limit the number of output tokens.
-We leave CoverUp’s target code segment size (see Section 3.1)
-at its default of 50 lines, and automatically repeat test executions
-to look for flaky tests (see Section 3.6).
-To place CoverUp on the same footing as CodaMosa, we first run it without any
-pre-existing test suite.
-We then run it two more times, allowing it to build upon the test suite from the
-previous runs.
-Since GPT-generated tests often assume the presence of certain Python modules,
-we configure CoverUp to automatically install any such modules if missing.
-4.2. Comparison to the previous state of the art
-To compare CoverUp to the previous state of the art, we first evaluate it on the CM
-suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
-Figure 10 shows, on the left, the line, branch and combined line and
-branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
-on the right, it shows the median of those measurements on a module-by-module basis.
-Additionally, Figure 12 plots the difference between the
-combined line and branch coverage obtained by CoverUp and CodaMosa on that suite,
-also on a module-by-module basis;
-green bars in the plot indicate modules where CoverUp achieved higher coverage.
-As the figures show, CoverUp achieves higher line, branch and combined
-line and branch coverage than both CodaMosa baselines, both measuring over the entire
-benchmark code base and on a per-module basis.
-Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage,
-49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage;
-on a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage,
-67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
-We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of
-CodaMosa (codex), with overall coverage measurements within 5% of each other.
+We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM, setting its “temperature” to zero and do not limit the number of output tokens.
+We leave CoverUp’s target code segment size (see Section 3.1) at its default of 50 lines and automatically repeat test executions to look for flaky tests (see Section 3.6).
+We first run it without any pre-existing test suite to place CoverUp on the same footing as CodaMosa.
+We then run it twice more, allowing it to build upon the test suite from the previous runs.
+Since GPT-generated tests often assume the presence of certain Python modules, we configure CoverUp to install any such missing modules automatically.
+4.2. [RQ1] Comparison to the previous state of the art
+To compare CoverUp to the previous state of the art, we first evaluate it on the CM suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
+Figure 12 shows, on the left, the line, branch, and combined line and branch coverage obtained by CoverUp and the baselines over the entire benchmark suite; on the right, it shows the median of those measurements on a module-by-module basis.
+Additionally, Figure 14 plots the difference between the combined line and branch coverage obtained by CoverUp and CodaMosa on that suite, also on a module-by-module basis; green bars in the plot indicate modules where CoverUp achieved higher coverage.
+Figure 12. [RQ1] CoverUp yields higher overall and median per-module coverage:
+Across the board, CoverUp yields higher coverage than CodaMosa, whether measured over the entire suite or on a module-by-module basis.
+Figure 13. [RQ1] CoverUp yields higher coverage than the state of the art:
+The graph shows the difference in (lines + branches) coverage between CoverUp and CodaMosa for the modules in the CM suite.
+Green highlights modules where CoverUp achieved a higher coverage.
+Figure 14.
+[RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage on all metrics, showing that its effectiveness is not limited to the modules selected for the CM suite.
+Figure 15. [RQ1] CoverUp yields higher overall and per-module coverage:CoverUp yields higher or equal coverage than MuTAP, whether measured over the entire suite or on a module-by-module basis.
+As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than both CodaMosa baselines, both measuring over the entire benchmark code base and on a per-module basis.
+Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
+On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
+These per-module improvements over CodaMosa (GPT-4o) and CodaMosa (Codex) are statistically significant: using paired permutation tests, we obtain a -value of for both, well below the standard threshold.
+We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of CodaMosa (codex), with overall coverage measurements within 5% of each other.
 As this difference shows, a newer model does not necessarily lead to higher performance.
-We next compare CoverUp to MuTAP, evaluating it on the MT suite.
-Figure 11 shows, on the left, the line, branch and combined line and
-branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
-on the right, it shows the median of those measurements on a module-by-module basis.
-As the figures show, CoverUp achieves higher line, branch and combined line and branch
-coverage than the MuTAP baselines when measuring over the entire benchmark,
-and greater or equal coverage when measuring on a per-module basis.
-Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage,
-89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage;
-on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics,
-except for MuTAP (gtp4o few-shot), which achieves extremely low coverage
-(17%, 0% and 11%).
-We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that
-its original few-shot prompt confuses GPT-4o, leading it to generate tests for
-the examples that make up the “few shot”, rather than for the function under test.
-When this happens, even if the resulting tests run to completion, they do not
-contribute to coverage.
-As the consistently high per-module median coverage values indicate, the MT suite does not
-contain particularly challenging code.
-In fact, the functions in the suite are entirely self contained, only rarely utilize external code,
-and when they do so, from standard libraries, and commonly include type annotations.
+As Figure 14 shows, CoverUp does not achieve higher coverage than
+CodaMosa for some modules.
+We examine CoverUp’s logs and observe that timeouts running tests for modules in the youtube_dl package constitute the largest single source of failure.
+While the authors of CodaMosa also report problems with timeouts for that package (Lemieux et al., 2023), it seems likely that CoverUp’s test execution timeout of one minute was set too low.
+In other cases, we observe that CoverUp’s static analysis cannot provide the LLM with correct context when the code under test contains conditional imports, such as those intended to accommodate package differences across Python versions.
+Additionally, while LLM test generations often include multiple test functions in the same response, CoverUp rejects them all if any of them fails; it is possible that CoverUp’s performance would have been higher if it were to evaluate each test function individually.
+Next, we evaluate the performance of CoverUp on code for which the original Pynguin performed well.
+Figure 14 shows the results.
+Given CoverUp’s 100% median per-module coverage and near 100% overall coverage results, we conclude that CoverUp also performs well on such code.
+We then compare CoverUp to MuTAP, evaluating it on the MT suite.
+Figure 15 shows, on the left, the line, branch, and combined line and branch coverage obtained by CoverUp and the baselines over the entire benchmark suite; on the right, it shows the median of those measurements on a module-by-module basis.
+As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than the MuTAP baselines when measuring over the entire benchmark and greater or equal coverage when measuring on a per-module basis.
+Across the entire benchmark suite, CoverUp achieves 90% (vs.
+82% to 16%) line coverage, 89% (vs.
+73% to 0%) branch coverage, and 90% (vs.
+79% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
+We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that its original few-shot prompt confuses GPT-4o: it generates tests for the test generation examples included in that prompt rather than for the function under test.
+When this happens, even if the resulting tests run to completion, they do not contribute to coverage.
+As the consistently high per-module median coverage values indicate, the MT suite does not contain particularly challenging code.
+In fact, the functions in the suite are entirely self-contained.
+They only rarely utilize external code; when they do so, they use it from standard libraries.
+Additionally, the functions commonly include type annotations.
 All of these characteristics greatly simplify the task of test generation.
-By contrast, the functions in the CM suite have numerous dependencies, often use external
-code and are generally unannotated.
-Figure 10. [RQ1] CoverUp yields higher overall and median per-module coverage:
-Across the board, CoverUp yields higher coverage than CodaMosa,
-whether measured over the entire suite or on a module-by-module basis.
-Figure 11. [RQ1] CoverUp yields higher overall and median per-module coverage:CoverUp yields higher or equal coverage than MuTAP,
-whether measured over the entire suite or on a module-by-module basis.
-Figure 12. [RQ1] CoverUp yields higher coverage than the state of the art:
-The graph shows the difference between the lines + branches coverage achieved by
-CoverUp and CodaMosa for the various benchmark modules in the CM benchmark suite,
-sorted by percentage of coverage increase.
-Green bars indicate modules where CoverUp achieved a higher coverage.
-Figure 13. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage
-on all metrics, showing that its effectiveness is not limited to
-the modules selected for the CM suite.
-Figure 14. [RQ3] Contribution of iterative coverage-guided approach:
-nearly 40% of the successes were obtained
-by iterating and providing coverage and error feedback,
-demonstrating its importance.
-We next evaluate the performance of CoverUp on code that the
-original Pynguin already performed well.
-Figure 14 shows the results.
-Given CoverUp’s 100% median per-module coverage and near 100% overall coverage
-results, we conclude that CoverUp also performs well on such code.
-[RQ1] Summary:CoverUp achieves significantly higher coverage than
-both CodaMosa and MuTAP, outperforming the state of the art.
-4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
-Given many LLMs’ near-human performance on various tasks, including software
-engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how
-much CoverUp contributes to performance.
-To address this question, we compare CoverUp to an ablated LLM-only
-version that utilizes a nearly identical prompt, but which lacks all other
-CoverUp functionality (coverage information, precomputed computed
-import statements, tool function for additional context
-and continued chats).
-As Figure 15 shows,
-CoverUp achieves substantially higher line, branch and combined line and branch coverage
-than CoverUp (ablated) both measuring over the entire benchmark code base
-and on a per-module basis.
-Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage,
-49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage;
-on a per-module basis, it achieves 82% (vs. 44%) line coverage,
-67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
+By contrast, the functions in the CM suite have numerous dependencies, often use external code, and are generally unannotated.
+[RQ1] Summary:CoverUp achieves significantly higher coverage than both CodaMosa and MuTAP, outperforming the state of the art.
+Overall Coverage
+Median Per-Module Coverage
+Test Generator
+Line
+Branch
+Line + Branch
+Line
+Branch
+Line + Branch
+CoverUp
+CodaMosa (codex)
+CodaMosa (gpt4o)
+CoverUp (ablated)
+CoverUpon PY suite
+Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and far outperforms itself when ablated to simply rely on LLM performance (top).
+Bottom line: as the results for the PY suite show, CoverUp’s performance is not limited to “challenging” code.
 Figure 16. [RQ2] CoverUp contributes significantly to overall performance:
-Across the board, CoverUp yields higher coverage in comparison to
-the ablated (LLM only) version, whether measured over the entire suite
-or on a module-by-module basis.
-[RQ2] Summary:CoverUp contributes significantly to overall performance;
-its performance is not just due to the LLM employed.
-4.4. How effective are CoverUp’s continued dialogues?
-To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
-identifying each successful test (i.e., which passes and improves coverage) that
-was generated immediately upon the first prompt or after continuing the chat with a second
-or third prompt
-(by default, and also in our evaluation, CoverUp continues the conversation for at most
-two additional prompts).
-Figure 14 shows the percentage of successes after
-the first, second and third prompts.
-As the figure shows, while the success rate decreases with each iteration, about
-40% of the successes were obtained through iterative refinement of the prompt,
-demonstrating its importance.
-[RQ3] Summary: continuing the chat contributes almost half its successes,
-demonstrating its effectiveness.
-5. Threats to Validity
-Benchmark selection
-We utilize the CM and PY benchmark suites to evaluate CoverUp’s
-performance on both challenging and less challenging code, as well
-as to facilitate including Codex-based results;
-we utilize the MT benchmark suite because it is supported by MuTAP
-and it, too, facilitates including Codex-based results.
-While so far our experience executing CoverUp to generate tests
-for other software has yielded similarly high coverage, selecting
-a different set of benchmarks could yield different results.
-Execution environment
-CodaMosa’s original evaluation environment failed to install a number of
-Python modules that are prerequisites for the applications used as benchmarks.
-In an effort to replicate the original conditions as closely as possible,
-we ignored these failures as well.
-It seems likely that both CodaMosa and CoverUp would be better able
-to generate tests if these modules were not missing.
-LLM model dependency
-CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo
-and GPT-4o models.
-While CoverUp’s test execution and coverage measurement portions
-are independent of the LLM
-and, as Section 4.3 shows, CoverUp significantly
-outperforms an ablated version of itself that just relies on the LLM’s capabilities,
-its ultimate performance naturally depends on the model’s ability to
-interpret its prompts and generate tests as requested.
-6. Conclusion
-This paper presents CoverUp, a novel approach for guiding
-LLM-based test generation with coverage analysis.
-We show that coupling coverage information with prompting, and iteratively
-refining prompts with updated coverage information to focus the LLM’s attention
-on improving uncovered code, is effective at generating test suites with
-higher coverage than past approaches.
-Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and
-far outperforms itself when ablated to simply rely on LLM performance.
+Across the board, CoverUp yields higher coverage in comparison to the ablated (LLM only) version, whether measured over the entire suite or on a module-by-module basis.
+4.3. [RQ2] How effective is CoverUp compared to simply prompting an LLM for tests?
+Given many LLMs’ near-human performance on various tasks, including software engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how much CoverUp contributes to performance.
+To address this question, we compare CoverUp to an ablated LLM-only version that utilizes a nearly identical prompt but which lacks all other CoverUp functionality (coverage information, computed import statements, get_info tool function and continued chats).
+Rather than ask for tests for the uncovered code, it asks instead for tests “that execute all lines and branches”.
+As Figure 16 shows, CoverUp achieves substantially higher line, branch, and combined line and branch coverage than CoverUp (ablated), both measured over the entire benchmark code base and on a per-module basis.
+Across the entire benchmark suite, it achieves 64% (vs.
+39%) line coverage, 49% (vs.
+19%) branch coverage, and 60% (vs.
+34%) line+branch coverage; on a per-module basis, it achieves 82% (vs.
+44%) line coverage, 67% (vs.
+7%) branch coverage, and 80% (vs.
+39%) line+branch coverage.
+[RQ2] Summary:CoverUp contributes significantly to overall performance; its performance is not just due to the LLM employed.
 Overall Coverage
@@
 Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
-metric
+4.4. [RQ3] How effective are CoverUp’s continued dialogues?
+To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite, identifying each successful test (i.e., which passes and improves coverage) that was generated immediately upon the first prompt or after continuing the chat with a second or third prompt (by default, and also in our evaluation, CoverUp continues the conversation for at most two additional prompts).
+We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.4% from the third.
+While the success rate decreases with each iteration, approximately 40% of successes were achieved through iterative refinement of the prompt, highlighting its importance.
+[RQ3] Summary: Continuing the chat contributes about 40% of its successes, demonstrating its effectiveness.
+4.5. [RQ4] How does the cost of running CoverUp compare to CodaMosa?
+Examining the logs from runs on the CM suite, we assemble Table 3, which shows the number of prompts, completions, tokens, running time, and approximate cost incurred.
+Test Generator
+Prompts
+P. Tokens
+Completions
+C. Tokens
+Time (h)
+Cost (US$)
+CoverUp
+28,690
+58,081,908
+28,654
+7,257,400
+4.0
+399
+CoverUp (ablated)
+9,224
+3,501,827
+9,188
+3,859,444
+1.7
+75
+CodaMosa (gpt4o)
+30,368
+39,508,559
+29,136
+4,504,329
+71.0
+265
+Table 3. [RQ4] Cost of running CoverUp and CodaMosa on the CM suite:CoverUp runs roughly 18 faster while achieving higher coverage
+than CodaMosa, but uses 48% more tokens.
+We observe that CoverUp runs roughly 18 times faster (4 vs. 71 hours) while achieving higher coverage than CodaMosa, but using 48% more tokens.
+Although this time comparison can provide insight into each method’s applicability for a given context, it is important to understand its limitations.
+CoverUp and CodaMosa utilize both local and cloud-based computing resources; we describe the local system in Section 4.1, but OpenAI does not disclose details of their computing environment, whose availability may vary over time.
+CoverUp’s implementation is both parallelized and asynchronous.
+Its running time is primarily gated by the rate limits imposed by OpenAI (see Section 3.7), their response times, and the time required to execute the generated tests.
+In contrast, CodaMosa is a sequential process that runs for 10 minutes per module.
+While this duration is configurable, we follow the setting used in the CodaMosa paper (Lemieux et al., 2023).
+[RQ4] Summary: As deployed, CoverUp runs 18 times faster while achieving higher coverage than CodaMosa, but using almost 50% more tokens.
+4.6. [RQ5] How important are CoverUp’s components to its performance?
+In broad terms, CoverUp contains components that perform three main functions: provide the LLM information on coverage, provide it with additional code context, and provide it with feedback on any errors.
+In this research question, we explore how ablating each one of these components affects CoverUp.
+Table 4 shows the specific elements disabled for each variant; disabling all three yields the same “CoverUp (ablated)” system explored already in RQ2.
+Component
+non-ablated
+ablated [RQ2]
+no coverage
+no code context
+no error fixing
+Indication of what coverage is missing
+Continued chat if generated tests do not improve coverage
+Generated import statements
+The get_info tool function
+Continued chat in case of errors
+Table 4.
+[RQ5] Elements in CoverUp ablations: The non-ablated CoverUp contains all of these elements, while CoverUp (ablated) from RQ2 contains none of them.
+To answer RQ5, we create three additional ablations.
+Figure 17. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
+Figure 18.
+[RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.
+We first evaluate these new ablations of CoverUp using suite CM; Figure 18 shows the resulting overall coverage for each ablation.
+We observe that all ablations negatively affect coverage, indicating their importance.
+Ablating error fixing has the most significant effect, losing between 14% and 37% on the various coverage metrics; ablating code context loses between 7% and 16%, and ablating coverage information up to 2%.
+In interpreting these results, it is important to realize that the components are not entirely independent.
+For example, the feedback from error fixing may allow the LLM to correct for insufficient code context, and through improved code context, errors resulting from incorrect assumptions may be avoided.
+To better understand the effect of ablating coverage information from the prompt, we count the test functions generated by the model and assess how many of these increase coverage for each successive run.
+As Figure 18 shows, as coverage increases, the coverage-ablated prompt results in fewer and fewer functions that increase coverage.
+In contrast, CoverUp’s prompt remains effective.
+This result shows that, by providing coverage information in the prompt, CoverUp focuses the LLM’s attention on uncovered code.
+Even though the coverage-ablated prompt only fell short in coverage by a small amount, it did so by generating over 50% more functions than the non-ablated prompt (11,222 vs. 7,366).
+If left in the test suite, these functions contribute to bloat, slowing the suite’s execution.
+But even if rejected after a coverage check, they add unnecessary tokens in LLM responses and thus increase cost.
+This analysis also suggests potential improvements for CoverUp.
+Since even the non-ablated prompt results in some test functions that do not increase coverage, CoverUp should measure the coverage generated by each function individually, rather than for the entire LLM response.
+Additionally, CoverUp could omit coverage information when prompting for tests for entirely uncovered functions, thereby reducing the cost of tagging every line in a code segment.
+We leave this to future work.
+Overall Coverage
+Median Per-Module Coverage
+Test Generator
 Line
 Branch
-l+b
-module median
-overall
-Table 3. [RQ1] CoverUp’s performance is not limited to “challenging” code: results for PY suite
-prompt
-contribution
-initial
-second
-third
-Table 4. [RQ3] Contribution of iterative coverage-guided approach to successes.
+Line + Branch
+Line
+Branch
+Line + Branch
+CoverUp (no coverage)
+CoverUp (no code context)
+CoverUp (no error fixing)
+CoverUp (ablated) [RQ2]
+Table 5. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
+First Run
+Second Run
+Third Run
+End
+Test Generator
+InitialCov.
+Cov. Incr.Functions
+InitialCov.
+Cov. Incr.Functions
+InitialCov.
+Cov. Incr.Functions
+Cov.
+TotalFunctions
+CoverUp
+7,366
+CoverUp (no coverage)
+11,222
+Table 6. [RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.
+[RQ5] Summary: Coverage-based prompting, code context, and error fixing are all important components of CoverUp; ablating any of them results in reduced coverage.
+5. Threats to Validity
+Benchmark selection
+We utilize the CM and PY benchmark suites to evaluate CoverUp’s performance on both challenging and less challenging code, as well as to facilitate including Codex-based results; we utilize the MT benchmark suite because MuTAP supports it, and it, too, facilitates including Codex-based results.
+While our experience executing CoverUp to generate tests for other software has yielded similarly high coverage, selecting a different set of benchmarks could produce different results.
+Execution environment
+CodaMosa’s original evaluation environment failed to install a number of Python modules that are prerequisites for the applications used as benchmarks.
+In an effort to replicate the original conditions as closely as possible, we ignored these failures as well.
+It seems likely that both CodaMosa and CoverUp would be better able to generate tests if these modules were not missing.
+LLM model dependency
+CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo and GPT-4o models.
+CoverUp’s approach is independent of the LLM and, as Section 4.3 shows, CoverUp significantly outperforms an ablated version of itself that just relies on the LLM’s capabilities.
+Its ultimate performance naturally depends on the model’s ability to interpret its prompts and generate tests as requested.
+6. Conclusion
+This paper introduces CoverUp, a novel approach to guide LLM-based test generation through coverage analysis.
+While modern LLMs achieve near-human performance on various tasks, simply prompting them is insufficient for generating high-coverage tests.
+We demonstrate that integrating coverage information into prompts directs the LLM’s attention to uncovered code and highlight the benefits of iterative refinement through error and coverage feedback.
+By combining these elements with code context, CoverUp produces test suites that achieve higher coverage than previous approaches.
+Alshahwan et al. (2024)
+Nadia Alshahwan, Jubin
+Chheda, Anastasia Finogenova, Beliz
+Gokkaya, Mark Harman, Inna Harper,
+Alexandru Marginean, Shubho Sengupta,
+and Eddy Wang. 2024.
+Automated Unit Test Improvement using Large
+Language Models at Meta. In Companion Proceedings
+of the 32nd ACM International Conference on the Foundations of Software
+Engineering, FSE 2024, Porto de Galinhas, Brazil, July 15-19, 2024,
+Marcelo d’Amorim (Ed.).
+ACM, 185–196.
+https://doi.org/10.1145/3663529.3663839
+Altmayer Pizzorno and Berger (2023)
+Juan Altmayer Pizzorno and
+Emery D. Berger. 2023.
+SlipCover: Near Zero-Overhead Code Coverage for
+Python. In Proceedings of the 32nd ACM SIGSOFT
+International Symposium on Software Testing and Analysis (Seattle, WA, USA)
+(ISSTA 2023). Association for
+Computing Machinery, New York, NY, USA,
+1195–1206.
+https://doi.org/10.1145/3597926.3598128
+arXiv:2305.02886
+Bareiß et al. (2022)
+Patrick Bareiß, Beatriz
+Souza, Marcelo d’Amorim, and Michael
+Pradel. 2022.
+Code Generation Tools (Almost) for Free? A Study of
+Few-Shot, Pre-Trained Language Models on Code.
+arXiv:2206.01335 [cs.SE]
+Boyapati et al. (2002)
+Chandrasekhar Boyapati,
+Sarfraz Khurshid, and Darko Marinov.
+2002.
+Korat: automated testing based on Java predicates.
+In Proceedings of the 2002 ACM SIGSOFT
+International Symposium on Software Testing and Analysis (Roma, Italy)
+(ISSTA ’02). Association for
+Computing Machinery, New York, NY, USA,
+123–133.
+https://doi.org/10.1145/566172.566191
+Bubeck et al. (2023)
+Sébastien Bubeck, Varun
+Chandrasekaran, Ronen Eldan, Johannes
+Gehrke, Eric Horvitz, Ece Kamar,
+Peter Lee, Yin Tat Lee,
+Yuanzhi Li, Scott Lundberg,
+Harsha Nori, Hamid Palangi,
+Marco Tulio Ribeiro, and Yi Zhang.
+2023.
+Sparks of Artificial General Intelligence: Early
+experiments with GPT-4.
+arXiv:2303.12712 [cs.CL]
+Chen et al. (2021)
+Mark Chen, Jerry Tworek,
+Heewoo Jun, Qiming Yuan,
+Henrique Ponde de Oliveira Pinto, Jared
+Kaplan, Harri Edwards, Yuri Burda,
+Nicholas Joseph, Greg Brockman,
+Alex Ray, Raul Puri,
+Gretchen Krueger, Michael Petrov,
+Heidy Khlaaf, Girish Sastry,
+Pamela Mishkin, Brooke Chan,
+Scott Gray, Nick Ryder,
+Mikhail Pavlov, Alethea Power,
+Lukasz Kaiser, Mohammad Bavarian,
+Clemens Winter, Philippe Tillet,
+Felipe Petroski Such, Dave Cummings,
+Matthias Plappert, Fotios Chantzis,
+Elizabeth Barnes, Ariel Herbert-Voss,
+William Hebgen Guss, Alex Nichol,
+Alex Paino, Nikolas Tezak,
+Jie Tang, Igor Babuschkin,
+Suchir Balaji, Shantanu Jain,
+William Saunders, Christopher Hesse,
+Andrew N. Carr, Jan Leike,
+Josh Achiam, Vedant Misra,
+Evan Morikawa, Alec Radford,
+Matthew Knight, Miles Brundage,
+Mira Murati, Katie Mayer,
+Peter Welinder, Bob McGrew,
+Dario Amodei, Sam McCandlish,
+Ilya Sutskever, and Wojciech Zaremba.
+2021.
+Evaluating Large Language Models Trained on Code.
+arXiv:2107.03374 [cs.LG]
+https://arxiv.org/abs/2107.03374
+Chen et al. (2024)
+Yinghao Chen, Zehao Hu,
+Chen Zhi, Junxiao Han,
+Shuiguang Deng, and Jianwei Yin.
+2024.
+ChatUniTest: A Framework for LLM-Based Test
+Generation. In Companion Proceedings of the 32nd
+ACM International Conference on the Foundations of Software Engineering,
+FSE 2024, Porto de Galinhas, Brazil, July 15-19, 2024,
+Marcelo d’Amorim (Ed.).
+ACM, 572–576.
+https://doi.org/10.1145/3663529.3663801
+Daka and Fraser (2014)
+Ermira Daka and Gordon
+Fraser. 2014.
+A Survey on Unit Testing Practices and Problems.
+In 25th IEEE International Symposium on Software
+Reliability Engineering, ISSRE 2014, Naples, Italy, November 3-6, 2014.
+IEEE Computer Society, 201–211.
+https://doi.org/10.1109/ISSRE.2014.11
+Dakhel et al. (2024)
+Arghavan Moradi Dakhel,
+Amin Nikanjam, Vahid Majdinasab,
+Foutse Khomh, and Michel C. Desmarais.
+2024.
+Effective test generation using pre-trained Large
+Language Models and mutation testing.
+Inf. Softw. Technol. 171
+(2024), 107468.
+https://doi.org/10.1016/J.INFSOF.2024.107468
+Fioraldi et al. (2023)
+Andrea Fioraldi,
+Alessandro Mantovani, Dominik Christian
+Maier, and Davide Balzarotti.
+2023.
+Dissecting American Fuzzy Lop: A FuzzBench
+Evaluation.
+ACM Trans. Softw. Eng. Methodol.
+32, 2 (2023),
+52:1–52:26.
+https://doi.org/10.1145/3580596
+Fraser and Arcuri (2011)
+Gordon Fraser and Andrea
+Arcuri. 2011.
+EvoSuite: automatic test suite generation for
+object-oriented software. In Proceedings of the
+19th ACM SIGSOFT Symposium and the 13th European Conference on Foundations of
+Software Engineering (Szeged, Hungary) (ESEC/FSE
+’11). Association for Computing Machinery,
+New York, NY, USA, 416–419.
+https://doi.org/10.1145/2025113.2025179
+Fraser and Arcuri (2014)
+Gordon Fraser and Andrea
+Arcuri. 2014.
 A Large-Scale Evaluation of Automated Unit Test

### v4

@@
 Compared to CodaMosa, a hybrid search/LLM-based test generator, CoverUp achieves a per-module median line+branch coverage of 80% (vs. 47%).
-Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 90% (vs. 77%).
+Compared to MuTAP, a mutation- and LLM-based test generator, CoverUp achieves an overall line+branch coverage of 89% (vs. 77%).
 We also demonstrate that CoverUp’s performance stems not only from the LLM used but from the combined effectiveness of its components.
 Test Generation, Regression Testing, Large Language Models, Code Coverage
-††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
+††copyright: acmlicensed††doi: 10.1145/3729398††journalyear: 2025††journal: PACMSE††journalvolume: 2††journalnumber: FSE††article: FSE128††publicationmonth: 7††ccs: Computing methodologies Artificial intelligence††ccs: Software and its engineering Software testing and debugging
 1. Introduction
@@
 45%).
-Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 90% (vs. 77%).
+Compared to MuTAP on one of the benchmark suites it supports, CoverUp again achieves greater or equal coverage on every metric, with an overall line+branch coverage of 89% (vs. 77%).
 This paper makes the following contributions:
@@
 CodaMosa (codex), the original version that uses the Codex LLM.
-Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp), available from https://github.com/microsoft/codamosa-dataset;
+Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp);
 CodaMosa (gpt4o), our adapted version that uses the same model as CoverUp, to help rule out differences in performance due to the use of different LLMs.
 MuTAP (codex few-shot), the original version that uses the Codex LLM, using a “few-shot” prompt, where two examples of test generation are included with the prompt.
-Since Codex is no longer available, we use original tests from MuTAP’s evaluation, available from https://github.com/ExpertiseModel/MuTAP;
+Since Codex is no longer available, we use original tests from MuTAP’s evaluation;
 MuTAP (codex zero-shot), also the original Codex version, but using
@@
 As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than both CodaMosa baselines, both measuring over the entire benchmark code base and on a per-module basis.
-Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
-On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
+Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 29%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
+On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 55% and 47%) line+branch coverage.
 These per-module improvements over CodaMosa (GPT-4o) and CodaMosa (Codex) are statistically significant: using paired permutation tests, we obtain a -value of for both, well below the standard threshold.
@@
 As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than the MuTAP baselines when measuring over the entire benchmark and greater or equal coverage when measuring on a per-module basis.
-Across the entire benchmark suite, CoverUp achieves 90% (vs.
-82% to 16%) line coverage, 89% (vs.
-73% to 0%) branch coverage, and 90% (vs.
-79% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
+Across the entire benchmark suite, CoverUp achieves 89% (vs.
+81% to 16%) line coverage, 89% (vs.
+73% to 0%) branch coverage, and 89% (vs.
+78% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
 We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that its original few-shot prompt confuses GPT-4o: it generates tests for the test generation examples included in that prompt rather than for the function under test.
@@
 [RQ1] Summary:CoverUp achieves significantly higher coverage than both CodaMosa and MuTAP, outperforming the state of the art.
+Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and far outperforms itself when ablated to simply rely on LLM performance (top).
+Bottom line: as the results for the PY suite show, CoverUp’s performance is not limited to “challenging” code.
 Overall Coverage
@@
 CoverUpon PY suite
-Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and far outperforms itself when ablated to simply rely on LLM performance (top).
-Bottom line: as the results for the PY suite show, CoverUp’s performance is not limited to “challenging” code.
 Figure 16. [RQ2] CoverUp contributes significantly to overall performance:
@@
 [RQ2] Summary:CoverUp contributes significantly to overall performance; its performance is not just due to the LLM employed.
+Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
 Overall Coverage
@@
 MuTAP (gpt4o few-shot)
-Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
 4.4. [RQ3] How effective are CoverUp’s continued dialogues?
 To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite, identifying each successful test (i.e., which passes and improves coverage) that was generated immediately upon the first prompt or after continuing the chat with a second or third prompt (by default, and also in our evaluation, CoverUp continues the conversation for at most two additional prompts).
-We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.4% from the third.
+We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.5% from the third.
 While the success rate decreases with each iteration, approximately 40% of successes were achieved through iterative refinement of the prompt, highlighting its importance.
@@
 Examining the logs from runs on the CM suite, we assemble Table 3, which shows the number of prompts, completions, tokens, running time, and approximate cost incurred.
+Table 3. [RQ4] Cost of running CoverUp and CodaMosa on the CM suite:CoverUp runs roughly 18 faster while achieving higher coverage
+than CodaMosa, but uses 48% more tokens.
 Test Generator
@@
 265
-Table 3. [RQ4] Cost of running CoverUp and CodaMosa on the CM suite:CoverUp runs roughly 18 faster while achieving higher coverage
-than CodaMosa, but uses 48% more tokens.
 We observe that CoverUp runs roughly 18 times faster (4 vs. 71 hours) while achieving higher coverage than CodaMosa, but using 48% more tokens.
@@
 Table 4 shows the specific elements disabled for each variant; disabling all three yields the same “CoverUp (ablated)” system explored already in RQ2.
+Table 4.
+[RQ5] Elements in CoverUp ablations: The non-ablated CoverUp contains all of these elements, while CoverUp (ablated) from RQ2 contains none of them.
+To answer RQ5, we create three additional ablations.
 Component
@@
 Continued chat in case of errors
-Table 4.
-[RQ5] Elements in CoverUp ablations: The non-ablated CoverUp contains all of these elements, while CoverUp (ablated) from RQ2 contains none of them.
-To answer RQ5, we create three additional ablations.
 Figure 17. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
@@
 But even if rejected after a coverage check, they add unnecessary tokens in LLM responses and thus increase cost.
-This analysis also suggests potential improvements for CoverUp.
-Since even the non-ablated prompt results in some test functions that do not increase coverage, CoverUp should measure the coverage generated by each function individually, rather than for the entire LLM response.
-Additionally, CoverUp could omit coverage information when prompting for tests for entirely uncovered functions, thereby reducing the cost of tagging every line in a code segment.
-We leave this to future work.
-Overall Coverage
-Median Per-Module Coverage
-Test Generator
-Line
-Branch
-Line + Branch
-Line
-Branch
-Line + Branch
-CoverUp (no coverage)
-CoverUp (no code context)
-CoverUp (no error fixing)
-CoverUp (ablated) [RQ2]
-Table 5. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
-First Run
-Second Run
-Third Run
-End
-Test Generator
-InitialCov.
-Cov. Incr.Functions
-InitialCov.
-Cov. Incr.Functions
-InitialCov.
-Cov. Incr.Functions
-Cov.
-TotalFunctions
-CoverUp
-7,366
-CoverUp (no coverage)
-11,222
-Table 6. [RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.
 [RQ5] Summary: Coverage-based prompting, code context, and error fixing are all important components of CoverUp; ablating any of them results in reduced coverage.
@@
 arXiv:2305.02886
+Altmayer Pizzorno and Berger (2025a)
+Juan Altmayer Pizzorno and
+Emery D. Berger. 2025a.
+plasma-umass/coverup: Automatically
+Generating Higher-Coverage Test Suites with AI.
+https://doi.org/10.5281/zenodo.15187805
+Altmayer Pizzorno and Berger (2025b)
+Juan Altmayer Pizzorno and
+Emery D. Berger. 2025b.
+plasma-umass/pytest-cleanslate: State
+pollution handling for CoverUp.
+https://doi.org/10.5281/zenodo.15187866
 Bareiß et al. (2022)
@@
 Few-Shot, Pre-Trained Language Models on Code.
+https://doi.org/10.48550/arXiv.2206.01335
 arXiv:2206.01335 [cs.SE]
@@
 experiments with GPT-4.
+https://doi.org/10.48550/arXiv.2303.12712
 arXiv:2303.12712 [cs.CL]
@@
 Evaluating Large Language Models Trained on Code.
+https://doi.org/10.48550/arXiv.2107.03374
 arXiv:2107.03374 [cs.LG]
-https://arxiv.org/abs/2107.03374
 Chen et al. (2024)
@@
 Current Large Language Models Versus Classical Optimizing Compilers.
+https://doi.org/10.48550/arXiv.2406.12146
 arXiv:2406.12146 [cs.AI]
-https://arxiv.org/abs/2406.12146
 Ryan et al. (2024)
@@
 (2020).
+https://doi.org/10.48550/arXiv.2009.05617
 arXiv:2009.05617
-https://arxiv.org/abs/2009.05617
 Turner (1986)
@@
 Engineering with ChatGPT.
+https://doi.org/10.48550/arXiv.2302.11382
 arXiv:2302.11382 [cs.SE]